### v2

# A DNN Biophysics Model with Topological and Electrostatic Features

Abstract
In this project, we provide a deep-learning neural network (DNN) based biophysics model to predict protein properties. The model uses multi-scale and uniform topological and electrostatic features generated using protein structural information and force field. The topological features are generated using the element specified persistent homology (ESPH) on a selection of heavy atoms or carbon atoms which contribute significantly in the protein biophysical properties. The electrostatic features are fast computed using a Cartesian treecode, modulated with treecode parameters for accuracy-cost tradeoff. These features are uniform in number for proteins with various sizes thus the broadly available protein structure database can be used to train the network. These features are also multi-scale thus the resolution and computational cost can be balanced by the users at the consideration of available computing powers. The machine learning simulations on two data sets with 4000+ and 17000+ protein structures showed the efficiency and fidelity of these features in representing the protein structure and force field for the predication of their biophysical properties such as electrostatic Coulombic energies and solvation energies. The feature generation algorithms have the potential as general tools in assisting machine learning based biophysical properties and function prediction for the broad biomolecules whose structures are obtained from both theoretical computing and experiments.
keywords:
Electrostatics; Poisson-Boltzmann; treecode; Deep Neural Network; Persistent Homology
1 Introduction
One of the overarching themes of biology is that structure determines function, which becomes more significant when we explore the biological world at the molecular level. With the advancement and availability of structure determination techniques such as X-ray crystallography [1, 2], NMR spectroscopy [3], cryo-electron microscopy [4], etc., there are rapidly growing 3D structures of protein, nucleic acids, and complex assemblies, etc. on publicly accessible repository such as Protein Data Bank (PDB) [5], GenBank [6], etc. These structure determination techniques have strengths and limitations that make them suitable for different types research studies. For examples, X-ray crystallography, due to the requirement of crystallization, is best for static high-resolution structures. NMR is valuable for studying dynamics in solution, particularly for the studies of smaller proteins and complexes. Cryo-electron microscopy which becomes more and more popular recently, excels in visualizing large, complex, and flexible biomolecular systems. The combination and integration of these methods due to their complementary natures help to reveal critical biomolecular functions embedded in the structures.
In addition to experiments, computing approaches are also available for the prediction of protein structures and properties. The main stream of Machine Learning based protein prediction focuses on using amino acid sequence to predict protein structure or structure-related properties, resulting in the trending protein language models (PLM) [7, 8], which is a transformer-based model inspired by Natural Language Processing (NLP) with wide range of applications such as protein pKa prediction [9], protein structure prediction [10], protein design [11], protein evolution [12], etc.
Recently, the appearance of AlphaFold developed by Google DeepMind revolutionarily reshaped the structures determination techniques. Based on known structures, AlphaFold learned to predict protein’s structures from sequence with very high accuracy as shown by their victory by a large margin in Critical Assessment of Techniques for Protein Structure Prediction (CASP 14-16), AlphaFold Protein Structures Database now has over 200 million protein structure predictions [13].
Both PLM and AlphaFold take protein sequences as inputs to predict structures or embeddings which keep important features, without touching the prediction of protein functions from protein structures. The machine learning methods potentially can revolutionize the traditional structure-based molecular modeling, which uses brownian dynamics, molecular dynamics, Monte Carlo simulations, etc. to predict protein properties and functions from protein structures. However, the development of ML methods in protein property predictions is hindered by the difficulty to represent protein structures and force fields using reliable features. To this end, the structure data, which is largely variant from protein to protein, needs to be converted to uniform size of features in representing structures to be loaded into the machine learning models such as regression models, random forest, deep neural network etc. Successful cases are the use of graph convolutional network (GCN) for predicting protein functions by features extracted from a protein language model and protein structure [14], protein-DNA binding specificity prediction using a wide range of features from geometry [15, 16], protein-ligand binding affinity prediction using topological features, [17, 18].
In this paper, we provide a novel way to generate electrostatic features. Combining electrostatic features with the topological features generated using a modification of the persistent homology approach used in [19], we represent the protein structures and associated force field in a uniform and multiscale fashion. With these features, we train a deep-learning neural network (DNN) to predict protein Coulombic energy and electrostatic solvation energy efficiently and accurately. Here the Solvation energies as labels are generated by solving the Poisson-Boltzmann (PB) model using a second-order Matched Interface and Boundary (MIB) method based PB solver [20]. There are recent research using Physics Informed Neural Network (PINN) to solve the Poisson-Boltzmann model [21, 22] while the present work is rather different as a Data-driven method with uniform and multi-scale features. In addition to these particular application in predicting Coulombic energy and electrostatic solvation energy, the feature generation algorithm we have developed potentially has wider application for other machine learning models as far as the protein structures and force field are required to be represented uniformly for proteins with various sizes.
This manuscript is constructed in the following way. After this section of introduction is the section of models and algorithms, followed by the section of results of simulation. The manuscript is ended with a conclusion section.
2 Models and Algorithms
We introduce the models and algorithms involved in this work. For protein’s structure-property relation we focus on the atomic description and the electrostatic interactions. The protein properties involved in this paper are Coulombic energy $E_{\text{coul}}$ and electrostatic solvation energy $E_{\text{solv}}$. For a protein, the Coulombic energy $E_{\text{coul}}$ is the energy it takes to displace atomic charges of the protein from infinity to their present positions (i.e. atomic centers from the PDB file) while the electrostatic solvation energy $E_{\text{solv}}$ is the energy it takes to solvate the protein into the solvent (e.g. water). Formal quantitative definition of these two protein properties will be given with mathematical models. For the protein-solvent interaction, we use the implicit solvent model in which water is represented as a continuum for the reduced computational cost. In a comparison perspective, we introduce the two most popular implicit solvation models, namely the Poisson-Boltzmann model and the Generalized Born’s model, which are used to generate the labels or features. Following that, we introduce the persistent homology and how we use it to produce the topological features. Finally we introduce how to use the Cartesian treecode, whose modification evolves into the algorithm to generate the electrostatic features.
2.1 The Poisson-Boltzmann model
Figure 1 depicts the popular implicit solvent models. In Fig.  1(a) a protein is represented by a collection of $N_{c}$ spherical atoms with centered partial charges. The molecular surface $\Gamma$ (also known as the solvent excluded surface) is defined by the trace of a water molecule represented by a red sphere rolling on contacting with the protein atoms. The Poisson-Boltzmann model is shown in Fig. 1(b), where the molecular surface $\Gamma$ divides the entire computational domain $\Omega$ into the protein domain $\Omega_{1}$ with dielectric constant $\varepsilon_{1}$ and atomic charges $q_{k}$ located at ${\bf x}_{k},k=1\,{:}\,N_{c}$, and the solvent domain $\Omega_{2}$ with dielectric constant $\varepsilon_{2}$ and dissolved salt ions  [23, 24]. Assuming a Boltzmann distribution for the ion concentration, and considering the case of two ion species with equal and opposite charges (e.g. Na+, Cl-), in the limit of weak electrostatic potential one obtains the linearized Poisson-Boltzmann (PB) model [25],
$\varepsilon_{1}\nabla^{2}\phi_{1}(\textbf{x})$ $=-\sum_{k=1}^{N_{c}}q_{k}\delta(\textbf{x}-\textbf{x}_{k}),\quad{\bf x}\in\Omega_{1},$ (1a)
$\varepsilon_{2}\nabla^{2}\phi_{2}(\textbf{x})$ $=\kappa^{2}\phi_{2}(\textbf{x}),\quad{\bf x}\in\Omega_{2},$ (1b)
$\phi_{1}(\textbf{x})$ $=\phi_{2}(\textbf{x}),\quad\varepsilon_{1}\frac{\partial\phi_{1}}{\partial n}({\bf x})=\varepsilon_{2}\frac{\partial\phi_{2}}{\partial n}({\bf x}),\quad{\bf x}\in\Gamma,$ (1c)
where $\kappa$ is the inverse Debye length measuring the salt concentration, and the potential satisfies a zero far-field boundary condition $\lim\limits_{|{\bf x}|\rightarrow\infty}\phi({\bf x})=0$.
[…]
(a)             (b)                 (c)
Figure 1: Implicit solvent models. (a) The Molecular Surface $\Gamma$: the trace of the solvent probe (shown in red) when it is rolled on contacting the spherical atoms of the protein (shown in green with a centered charge); (b) The Poisson–Boltzmann Model: two domains $\Omega_{1}$ (dielectric constant $\epsilon_{1}$ with partial charges as weighted summation of delta functions) and $\Omega_{2}$ (dielectric constant $\epsilon_{2}$ with mobile ions modeled by Boltzmann distribution) are separated by the molecular surface $\Gamma$; (3) The Generalized Born Model: the protein is represented as a collection of $N_{c}$ spherical atoms centered at ${\bf x}_{k}$ with charge $q_{k}$ and Born Radius $R_{k}$ (only the $k$th atom is shown).
The PB model governs the electrostatic potential $\phi$ in the entire space. Theoretically after $\phi$ is obtained, its gradient will produce electrostatic field while its integral will generate potential energy. However, there are many challenging issues on properly obtaining the field and energy (e.g. definition of field on molecular surface $\Gamma$ [26]). Our attention for this project is on the energy side as described below. The electrostatic free energy is given as
$E_{\text{free}}=\frac{1}{2}\int_{\Omega}\rho({\bf x})\phi({\bf x})d{\bf x}=\frac{1}{2}\sum\limits_{k=1}^{N_{c}}q_{k}\phi({\bf x}_{k})=\frac{1}{2}\sum\limits_{k=1}^{N_{c}}q_{k}(\phi_{\text{reac}}({\bf x}_{k})+\phi_{\text{coul}}({\bf x}_{k}))=E_{\text{solv}}+E_{\text{coul}}$ (2)
where $\rho({\bf x})=\sum\limits_{k=1}^{N_{c}}q_{k}\delta(\textbf{x}-\textbf{x}_{k})$ is the charge density as a sum of partial charges weighted delta function and the $E_{\text{solv}}=\frac{1}{2}\sum\limits_{k=1}^{N_{c}}q_{k}\phi_{\text{reac}}({\bf x}_{k})$ term is the electrostatic solvation energy. The $\phi_{\text{reac}}({\bf x}_{k})$ is the reaction potential at ${\bf x}_{k}$ as the remaining component when Coulomb potential $\phi_{\text{coul}}({\bf x}_{k})=\sum\limits_{j=1,j\neq k}^{N_{c}}\frac{q_{j}}{\epsilon_{1}|{\bf x}_{k}-{\bf x}_{j}|}$ is taken away from the total electrostatic potential $\phi({\bf x}_{k})$.
Solving the PB model numerically by grid-based methods is challenging because (1) the protein is represented by singular point charges, (2) the molecular surface is geometrically complex, (3) the dielectric constant is discontinuous across the surface, (4) the domain is unbounded. To overcome these numerical difficulties, several finite difference interface methods have been developed. By assuming that the interface is aligned with a mesh line, a jump condition capture scheme has been developed in [27]. Based on a Cartesian grid, the immersed interface method (IIM) [28] has been applied to solve the PB equation [29, 30], in which the jump conditions can be rigorously enforced based on Taylor expansions. For the purpose of dealing with arbitrarily shaped dielectric interfaces based on a simple Cartesian grid, a matched interface and boundary (MIB) PB solver [31, 32, 33, 20, 34] has been developed through rigorous treatments of geometrical and charge singularities. Alternatively, boundary element methods (BEM) for the PB model were developed later [35, 36, 37, 38, 39, 40, 41, 42] with several inherent advantages, (1) only the molecular surface is discretized rather than the entire solute/solvent volume, (2) the atomic charges are treated analytically, (3) the interface conditions are accurately enforced, (4) the far-field boundary condition is imposed analytically, (5) the matrix-vector product in iterative methods can be accelerated by fast algorithms [43, 44, 45, 46]. In this paper, we computed the solvation energies using the MIBPB solver [20] since accuracy is of critical importance for the generation of labels ($E_{\text{solv}}$) that will be used to train the neural network.
2.2 Generalized Born’s Model
If ion effects are not considered, the solvation energy and the reaction potential can also be efficiently computed by using the generalized Born’s (GB) model, an approximation to the Poisson’s model. The GB model has significantly reduced computational cost as opposed to the PB or Poisson models, whose solutions require solving a 3D partial differential equation. Although we did not use the GB model to produce any numerical results for this project, we describe the GB model for its relationship with the PB model and its potential usage in electrostatic interactions. In fact, the GB model can calculate the electrostatic solvation energy which can be used either as the label or as a feature [47, 48] in our machine learning framework. It can also produce the reaction potentials at the atomic center, which can be used to generate the second level of electrostatic features as shown in Fig. 3, which will be explained within that content.
To derive the GB model, we first adopt some classic results for electrostatics from Jackson’s textbook [49] (sec. 1.11). Using the analogue from discrete point charge distribution for continuous charge density distribution, the electrostatic free energy of a charged system with density distribution $\rho({\bf r})$ takes the form
$E_{\text{free}}=\frac{1}{2}\int\rho({\bf r})\phi({\bf r})d{\bf r}=\frac{1}{8\pi}\int{\bf E}({\bf r})\cdot{\bf D}({\bf r})d{\bf r}$
by using the Poisson equation $\nabla^{2}\phi=\rho({\bf r})/\epsilon$, the electric field ${\bf E}=-\nabla\phi$, and electric displacement ${\bf D}=\epsilon{\bf E}$.
Now consider assembling a charge $q_{i}$ to the center of a sphere of radius $a_{i}$ and the center is located at the origin. The sphere separates the domain with $\epsilon_{1}$ for $r<a_{i}$ and $\epsilon_{2}$ for $r>a_{i}$. The assembly takes the energy
$E_{\text{free,}i}=\frac{1}{8\pi}\int\frac{{\bf D}\cdot{\bf D}}{\epsilon}d{\bf r}\approx\frac{1}{8\pi}\int_{r<a_{i}}\frac{q_{i}^{2}}{r^{4}\epsilon_{1}}d{\bf r}+\frac{1}{8\pi}\int_{r>a_{i}}\frac{q_{i}^{2}}{r^{4}\epsilon_{2}}d{\bf r}$ (3)
where the Coulomb field approximation ${\bf D}=\displaystyle\frac{q_{i}{\bf r}}{r^{3}}$ is used.
From a slightly different perspective as in Eq. (2), the electrostatic solvation energy is the energy difference when $\epsilon$ for $r>a_{i}$ changes from $\epsilon_{1}$ (unsolvated) to $\epsilon_{2}$ (solvated) given as
$E_{\text{solv},i}=\frac{1}{8\pi}(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}})\int_{r>a_{i}}\frac{q_{i}^{2}}{r^{4}}d{\bf r}=(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}})\frac{q_{i}^{2}}{2a_{i}}.$ (4)
This result is consistent with the Poisson-Boltzmann model of a solvated spherical cavity with centered charges as summarized in [50], which is a special case from Kirkwood’s derivation of a series of spherical harmonics for a spherical cavity containing arbitrary multiple charges [51]. If we treat the molecule as a collection of spherical atoms, using Eq. (4) the total electrostatic solvation energy is
$E_{\text{solv}}=\frac{1}{2}\left(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}}\right)\left(\sum_{i=1}^{N}\frac{q_{i}^{2}}{a_{i}}+\sum_{j\neq i}^{N}\frac{q_{i}q_{j}}{r_{ij}}\right)\approx\frac{1}{2}(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}})\sum_{i=1}^{N}\sum_{j=1}^{N}\frac{q_{i}q_{j}}{f_{ij}^{\text{GB}}}$ (5)
where $r_{ij}$ is the distance between the atomic centers of atoms $i$ and $j$. Here the first equation of $E_{\text{solv}}$ has two terms: a sum of individual Born’s terms and pairwise Coulombic terms. Note the actual molecule is better represented by a dielectric interface (e.g. solvent excluded surface (SES)) which separates inside domain $\Omega_{1}$ and outside domain $\Omega_{2}$. The charge $q_{i}$ is located at the center ${\bf r}_{i}$ of a sphere with radius $a_{i}$. We assume $\Omega_{1}$ contains all these spheres. The model using dielectric interface requires a step further as been approximated in the second equation in Eq. (5).
Here $f_{ij}^{\text{GB}}$ is the effective Born radii ($i=j$) or effective interaction distance ($i\neq j$). Assuming the effective Born Radii $R_{i}$’s are obtained, a popular estimation of $f_{ij}^{\text{GB}}$ is:
$f_{ij}^{\text{GB}}(r_{ij})=(r_{ij}^{2}+R_{i}R_{j}e^{\frac{-r_{ij}^{2}}{4R_{i}R_{j}}})^{\frac{1}{2}}$
Note $R_{i}$ depends not only on $a_{i}$, but also on radii and relative positions of all other atoms. The estimation of effective Born radii is an active research area [52]. From Poisson-Boltzmann theory, the perfect Born radii is given as
$R_{i}=\frac{1}{2}(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}})\frac{q_{i}}{E^{\text{PB}}_{\text{solv},i}}$
with $E^{\text{PB}}_{\text{solv},i}$ as the electrostatic solvation energy from the PB model with all except $i$th charge muted. This PB-GB relationship can be seen in Fig. 1(c), in which the electrostatic solvation energy of the sphere equals that of the molecule, both with a charge at the atomic center. There are many approaches to approximate $R_{i}$ and a typical one is the volume integration via quadrature as
$R_{i}^{-1}=\frac{1}{4\pi}\int_{\Omega_{2}}\frac{q_{i}^{2}}{r^{4}}d{\bf r}=\frac{q_{i}^{2}}{a_{i}}-\frac{1}{4\pi}\int_{\Omega_{1}/B({\bf r}_{i},a_{i})}\frac{q_{i}^{2}}{r^{4}}d{\bf r}.$
2.3 Topological Features
This section involves many fundamental concepts in Algebraic Topology. We use italic fonts to emphasize important terms and use bold fonts to indicate their sequential definition if there are any. The fundamental task of topological data analysis is to extract topological invariants as the intrinsic features of the underlying space. For the prediction of protein properties, we expect the topological invariants such as independent components, rings, cavities, etc. carry useful information which cannot be discovered from geometric observation and measurement. This is realized by the aid of homology, a mathematical framework that assigns algebraic objects (groups) to a topological space to measure its shape, especially its holes in various dimensions. Below we briefly introduce the simplicial homology and the persistent homology, and their associated terms and definitions. Following that, we explain how these homologies and their computations with protein structures are used to generate topological features.
2.3.1 Simplicial Homology
Topological invariants in a discrete data set (e.g. the collection of atomic positions of a protein) can be studied using simplicial homology which uses a specific rule (e.g. the choice of proximity parameter $\epsilon$ and an associated procedure to determine the edges connecting the vertices) to identify simplicial complexes (will be defined later) from simplexes. Here the simplex represents the simplest possible polytope in any given dimension like point, line segment, triangle, tetrahedron, etc. Formally, A $k$-simplex is the convex hull of $k+1$ affinely independent points $v_{0},v_{1},\dots,v_{k}$:
$\sigma^{k}=\left\{\sum_{i=0}^{k}\lambda_{i}v_{i}\;\middle|\;\lambda_{i}\geq 0,\sum\lambda_{i}=1\right\}$
Each subset of $m+1$ of these vertices forms an $m$-face of the simplex, which is itself an $m$-simplex.
The boundary of a $k$-simplex $v_{0},\dots,v_{k}]$ is:
$\partial_{k}[v_{0},v_{1},\dots,v_{k}]=\sum_{i=0}^{k}(-1)^{i}[v_{0},\dots,\hat{v}_{i},\dots,v_{k}]$
where $\partial_{k}$ is the boundary operator from $k$-chains to (k1)(k-1)-chains and $\hat{v}_{i}$ means vertex $v_{i}$ is omitted. Here a $k$-chain is a formal linear combination of k-dimensional simplexes in a simplicial complex.
With the concepts of simplex, face, boundary, chain, etc., we can formally define that a simplicial complex $K$ is a collection of simplices such that:
  1. 1.
    Every face of a simplex in $K$ is also in $K$.
  2. 2.
    The intersection of any two simplices is either empty or a common face.
Thus, each simplicial complex corresponds to a topological space by “filling in” the simplexes and gluing them together along shared faces.
2.3.2 Persistent Homology
However, the simplicial homology of a complex associated to a point set at a particular proximity parameter $\epsilon$ is insufficient to characterize the signal or noise. This calls for the ideas of filtration and persistent homology, which can identify and connect complexes at different level of complexity and record the appearance and disappearance of the homology group. The importance of persistence is that given a parameterized family of spaces, topological features which persist over a significant parameter ranges are considered as signal while those with short-lived features are treated as noise [53]. To this end, we introduce the concept of filtration, which is a sequence of nested simplicial complexes:
$\emptyset=K_{0}\subseteq K_{1}\subseteq\cdots\subseteq K_{n}=K.$
Each $K_{i}$ is a simplicial complex, and the sequence grows over time or scale. Instead of computing just the $k$-homology group $H_{k}(K)$, which is an algebraic object capturing the $k$-dimensional holes in a topological space represented by a simplicial complex $K$, we compute $H_{k}(K_{i})$ for each $i$ and track features as the birth, death, and persistence of a homology class. Let $C_{k}(K)$ be the group of $k$-chains, and let
$\cdots\xrightarrow{\partial_{k+1}}C_{k}(K)\xrightarrow{\partial_{k}}C_{k-1}(K)\xrightarrow{\partial_{k-1}}\cdots$
be the sequence of boundary maps, the $k$-homology group of $K$ is defined as:
$H_{k}(K)=\frac{\ker(\partial_{k})}{\operatorname{im}(\partial_{k+1})}$
where
  • 1.
    $\ker(\partial_{k})$ is the group of $k$-cycles: $Z_{k}=\{c\in C_{k}\mid\partial_{k}(c)=0\}$
  • 2.
    $\operatorname{im}(\partial_{k+1})$ is the group of $k$-boundaries: $B_{k}=\{\partial_{k+1}(c)\mid c\in C_{k+1}\}$
Hence, the $k$-th homology group is the group of cycles modulo boundaries:
$H_{k}=Z_{k}/B_{k}$
Taking $H_{1}$ as an example, $Z_{1}=\ker(\partial_{1})$ collects 1-chains (formal sum of edges) without boundary points thus general loops, while $B_{1}=\operatorname{im}(\partial_{2})$ collects boundaries of 2-chains (a collection of triangles) that bound a filled-in region thus trivial holes, $H_{1}=Z_{1}/B_{1}$ is then the collection of actual holes in the space.
2.3.3 Feature Generation from Barcode
To make the topological features uniform and physically informed, we choose to use Element Specific Persistent Homology (ESPH) to extract topological features at different level of complexity. ESPH has been used for successfully predicting protein-ligand binding [19]. The main idea is to select atom elements, which form the cloud of points, using target problem related physical, chemical, and biological information. In this project, since solvation energy and Coulomb energy are highly related to the amino acid chains, we use two collections of cloud of points. One is the collection of all Carbon atoms, which form the backbone of amino acids. The other is the collection of all heavy atoms {C, N, O, S}, whose locations and interconnections are also critical in protein structures. Once the collections of cloud of points are determined and selected, the barcode [53] carrying persisting homology information of the cloud of points are produced. Here the barcode is a graphical representation of persistent homology as a collection of horizontal line segments in a plane whose horizontal axis corresponds to the proximity parameter and whose vertical axis represents an (arbitrary) ordering of homology generators such as $H_{0}$, $H_{1}$, $H_{2}$, etc. We use the software GUDHI [54] to implement this task. The produced barcode is then used by the algorithm below to generate the vector of topological features.
Figure 2: The barcode generated by the collection of cloud points using heavy atoms in proteins (a) 1a1e and (b) 1pb9.
Figure 2 gives the barcode for protein 1a1e, the C-SRC Tyrosine Kinase, and protein 1pb9, the N-methyl-D aspartate receptor, using heavy atoms. The persistent homology analysis conducted in this study characterizes key biophysical phenomena, including intermolecular interactions, solvation effects, and hydrophobicity. Topological features are extracted from Betti-1 and Betti-2 barcodes, where the birth, death, and persistence values are computed based on the spatial distribution of carbon atoms and all heavy atoms within the protein structure, resulting in a total of 12 feature channels. Features derived from all heavy atoms represent the geometric interactions between specific types of atoms in the protein, whereas those computed from carbon atoms primarily reflect hydrophobic interactions and indirectly capture solvation effects. The analysis is performed over a distance scale of 0 to 50Å, which is discretized into uniform bins of 0.50Å to ensure fine-grained resolution of topological patterns. The topological information included in the barcode can be converted to features with the following approach [19].
We define the collection of barcodes as $\mathbb{B}(\alpha,\mathcal{C},\mathcal{D})$ with the following specifications.
$\alpha$ : atom labels (a specific way to obtain a collection of atoms i.e., all heavy atoms in protein, heavy atoms close to the ligand, atoms near the mutated residue, etc.)
$\mathcal{C}$ : type of simplicial complex (i.e., Ribs or Cech)
$\mathcal{D}$ : dimension (i.e., Betti-0, Betti-1, etc.)
Using the collection, the structured vectors ${\bf V}^{b}$, ${\bf V}^{d}$, and ${\bf V}^{p}$ can be constructed to respectively describe the birth, death, and persistent patterns of the barcodes in various spatial dimensions. Practically, the filtration interval $[0,L]$ is divided into $n$ equal length subintervals and the patterns are characterized on each subinterval. The description vectors using their $i$th component are defined as:
[…]
2.4 Electrostatic Features
2.4.1 Introduction
In most of the molecular simulations, including Monte Carlo simulation, Brownian Dynamics, Molecular Dynamics, etc., the electrostatic interactions are characterized by the partial charges assigned at the atomic centers. These partial charges are assigned according to the force fields, which are a set of mathematical equations and parameters used to calculate the potential energy of a molecule based on its atomic structure and interactions. The force fields are determined by experiment or quantum chemistry.
Figure 3: An 2-d illustration for 3-d uniform and multi-scale electrostatic features for a protein (purple dashed line) with charges (shown as black dots); left: charges $q$ and reaction potential $\phi_{\text{reac}}$ are redistributed as point-multipoles (shown as explosion symbols) using Cartesian treecode [45] or FMM [46] at the centers of the cluster at different levels (level 0: black; level 1: red; level 2: green); right: charges $q$ and reaction potential $\phi_{\text{reac}}$ are redistributed at Chebyshev nodes (shown as square: level 0: black; level 1: red; level 2: green; only one cluster is shown in each level) using Barycentric Lagrangian treecode [56].
The algorithm for obtaining the mulit-scale, physics-informed, uniform electrostatic features can be explained using the left of Fig. 3. The electrostatic profile of the protein is determined by the partial charges $q_{n}({\bf r}_{n})$ or reaction potential $\phi_{\text{reac}}({\bf r}_{n})$ at the atomic centers ${\bf r}_{n}$’s as illustrated by the black dots for $n=1,2,\cdots,N_{c}$ with $N_{c}$ as the number of atoms of the protein. Partial charges are the source of electrostatic interactions while reaction potentials are the outcomes of electrostatic interactions for the solvated potein, which includes the important solute-solvent interactions. Consider the variation of $N_{c}$, the number of atoms, from protein to protein, it is difficult to directly use $q_{n}({\bf r}_{n})$ and/or $\phi_{\text{reac}}({\bf r}_{n})$ as features for machine learning. Our strategy to alleviate this pain is to use point-multipoles as illustrated by the explosion symbols at the cluster centers to carry redistributed charges or reaction potentials in terms of moments of the multipole expansion. The number of cluster centers are uniform for proteins with various numbers of atoms. It is also multiscale as being determined by the combination of levels of the tree $L$ and number of terms of the multipole expansion $p$. The users specify $L$ and $p$ as needed.
In the left of Fig. 3. the $N_{d}$ point multipoles (explosion symbols) are organized in a hierarchical way such that different level of details are given by different number of symbols at different levels. For the $m$th multipole, $m=1,\cdots,N_{d}$, using a 3-d index ${\bf k}=(k_{1},k_{2},k_{3})$, $|{\bf k}|=k_{1}+k_{2}+k_{3}$, in a cluster $c$ with cluster center ${\bf x}_{c}$, the moments are defined as
$\mathcal{M}_{c}^{\bf k}=\sum\limits_{{\bf x}_{j}\in c}q_{j}({\bf x}_{j}-{\bf x}_{c})^{\bf k}.$ (7)
Note when $|{\bf k}|=0$, ${\bf k}=(0,0,0)$; when $|{\bf k}|=1$, ${\bf k}\in\{(1,0,0),(0,1,0),(1,0,0)\}$; when $|{\bf k}|=2$, ${\bf k}\in\{(1,1,0),(1,0,1),(0,1,1),(2,0,0),(0,2,0),(0,0,2)\}$, etc.
If the Cartesian treecode algorithm [45] is used, the computational cost is $O(N_{c}\log N_{c})$ for obtaining all the moments as detailed in [45], where $N_{c}$ is number of the particles/charges and $\log N_{c}$ is the level of trees. In fact, the computational cost can be as low as $O(N_{c})$ using the strategies in [46] such that the moments at the finest cluster are computed first and a M2M (moments to moments) transformation can be used to efficiently compute moments at any desired level from the bottom to the top. Our simulations so far showed that the $O(N_{c}\log N_{c})$ cost is sufficiently fast for moments computation since $N_{c}$ is within the scale of tens of thousands even for large proteins. These moments are intrinsic properties of the cluster thus can serve as features for the protein, which carries simplified but important structural and charge information.
The number of features up to level $L$ cluster (the largest box containing all particles is named as the level 0 cluster and each subdivision increases the level by 1) is
$N_{f}(p,L)=N_{p}(1+8+8^{2}+\cdots+8^{L})=N_{p}\frac{8^{L+1}-1}{7}$
where $N_{p}$ is the number of terms in multipole expansion with $N_{p}=1,4,10,20,35,56\cdots$ as the sequence of tetrahedral numbers when $p$th order multipole is used.
Alternatively, the Barycentric treecode approach as shown in the right picture of Fig. 3 can be used to generate the electrostatic features located the Chebyshev points (the squares) with modified weights as feature values. The details can be found in [56]. For this approach, the computational cost is $O(N_{c}\log N_{c})$, and the number of terms in the similar sense of $N_{p}$ is $O(c_{p}^{3})$ where $c_{p}$ is the number of Chebyshev nodes in one direction. In this project, we focus on the Cartesian treecode approach, while the Barycentric treecode approach will be investigated in the future.
2.4.2 Implementation details
We provide the step-by-step procedure by which the electrostatic features are generated.
Step 1. The protein structure (e.g. the PDB file) is obtained from publicly available repo e.g. the protein data bank [5]. The protein is represented by its atomic locations and partial charges: $q_{i}({\bf r}_{i})$ for $i=1,\cdots,N_{c}$, e.g. the PQR file using the PDB2PQR software [57].
Step 2. Two parameters $p$ and $L$ are given, where $p$ is the order in the Cartesian Taylor expansion and $L$ is the number of levels. With given $p$, we have the number of terms given as $N_{p}=\displaystyle{(p+1)(p+2)(p+3)}/{6}$. The number of features up to level $L$ cluster is
$N_{f}(p,L)=\frac{p(p+1)(p+2)(8^{L+1}-1)}{42}.$
These $N_{f}$ numbers are ordered by level from 0 to $L$ and the coordinates of cluster centers in each level. This determines the dimension of our final feature vector $F\in\mathbb{R}^{N_{f}}$.
The following table gives how $N_{f}$ varies with the change of $p$ and $L$:
$p\LL$ 0 1 2 3 4
0 1 9 73 585 4681
1 4 36 292 2340 18724
2 10 90 730 5850 46810
3 20 180 1460 11700 93620
4 35 315 2555 20475 163835
Step 3. The Cartesian treecode [45] will be called and the moments for the multipoles will be output as the electrostatic features. These features can be used as part of the inputs for machine learning algorithms, which will be shown the next.
2.5 A DNN based machine learning model with topological and electrostatic features
Figure 4: The DNN based Machine Learning model which uses protein structural data, force field, and known protein properties to mathematically generate electrostatic, topological, algebraic, geometric, and features to train a learned model and then use the learned model to predict unknown properties for protein with available structures.
The main goal of this project is to discover the hidden and useful information embedded in the protein structural data from the protein data bank in a simple and abstract way. The protein structure data and force field include bonded and non-bonded interactions with which the molecular simulation can be performed. We categorize these information into electrostatic, topological, algebraic, and geometric features. These four aspects are all important to the machine learning model. In this project, we put our focus on the electrostatic and topological features. The DNN based Machine Learning model is shown in Fig. 4, which uses the available protein structural data [58], force field such as AMBER [59], CHARMM[60], AMOEBA [61], etc. and known protein properties repositories such as PDBbind Database [62], Protein pKa Database [63], etc to mathematically generate features to train a learned model and then use the learned model to predict unknown properties for protein with available structures. In this project, we focus on the effectiveness of topological and electrostatic features using computed Coulombic energies and solvation free energies as labels.
3 Results
3.1 Specification of the machine learning model
We provide more details of the machine learning model with focus on the neural network and its related specification. The predictive model developed in this study is designed to predict solvation and Coulomb energies based on a combination of topological and electrostatic features. Prior to model training, outliers are identified and removed using the interquartile range (IQR) method. All input features and target labels are standardized to have zero mean and unit variance using the StandardScaler() utility from sklearn.preprocessing. The inverse transformations are applied to the output, returning values with physical meanings e.g. energies in kcal/mol during post-processing for evaluation purposes. The network architecture of the model as shown in Fig. 5 is implemented using the Keras Sequential API with two distinct branches. The first branch as color-coded in blue is a one-dimensional convolutional neural network (CNN) processing the topological features. The CNN branch has three convolution-pooling layers with (128, 64, 64) filters and ($\tanh$, relu, $\tanh$) activation function respectively. The 3 by 3 kernels are used. This CNN branch is responsible for capturing hierarchical and spatially correlated patterns in the topological data. The output of the CNN is flattened and passed through a subsequent Dense layer of 64 units. The other branch is the electrostatic features into two fully connected hidden dense layers with (128, 64) units respectively as shown in orange. After independent processing, the outputs from both branches are concatenated and passed through additional three hidden dense layers with (32, 16, 8) units as shown in purple to produce the final regression output of energy as shown in red.
Figure 5: The architecture of the deep neural network with two branches taking in topological features and electrostatic features. The outputs of the two branches are concatenated into one deep neural network.
To enhance generalization and mitigate overfitting, batch normalization is applied after intermediate layers, and dropout is employed to randomly deactivate neurons during training. The model is trained for 500 epochs using the Adam optimizer with a learning rate of 0.0001 and a batch size of 32. The mean squared error (MSE) is used as a loss function. Model architecture and hyperparameters are selected through empirical tuning and informed by previous studies employing Hyperopt [19, 64]. The dataset as described below was split randomly into 80% for training and 20% for testing, and the model was trained with 5-fold cross-validation as well.
3.2 Data
The datasets are sourced from the PDBbind database [65]. In particular, we categorize used data into two datasets as described below. The Amber ff14SB general force field is applied for the atomic van der Waals radii and partial charges when the PDB2PQR [57] software is used.
Dataset 1: These data are from the refined set of the v2018 release [65]. An interquartile range method is used to remove outliers based on $E_{\text{coul}}$, resulting in 4,658 complexes used in our simulation. We also we computed the electrostatic solvation free energies $E_{\text{solv}}$ of the proteins, as part of the 4,658 protein-ligand complexes. All the $E_{\text{solv}}$ values are computed using the 2nd order accurate Matched Interface and Boundary Poisson-Boltzmann (MIBPB) solver [20] at mesh size h=0.2h=0.2Å.
Dataset 2: These data are from the the v2020 release, which has a general set with 14,127 complexes and refined set with 5,316 complexes. We computed the Coulomb energies $E_{\text{coul}}$ of the all proteins using the Treecode [45]. An interquartile range method is used to remove outliers based on $E_{\text{coul}}$, resulting in 17,007 complexes used in our simulation.
3.3 Research Purposes and Metrics
We here provide some simulation results to demonstrate the performance of our model. These results serve three research purposes.
Purpose 1: Training on larger data set receives better performance.
Purpose 2: When electrostatic features are considered, performance of model improves when when treecode level $L$, Taylor expansion order $p$, and their associated number of features $N_{f}$ increases.
Purpose 3: Using both electrostatic features and topological features outperforms using only one kind of these features.
Metrics: We use three common metrics to measure the performance of the model as defined below. In all these metrics, we use $y_{i}$ to denote actual value, $\hat{y}_{i}$ for predicted value, $\bar{y}$ for mean of actual values, $\bar{\hat{y}}$ for mean of predicted values.
MAPE: The mean average percentage errors, which measures prediction accuracy as a percentage. Not suitable if the actual values contain zero or values close to zero.
$\text{MAPE}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|\times 100\%$
MSE: The mean-squared error, which measures the average absolute difference between predicted and actual values.
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$
$R^{2}$, which indicates the proportion of variance in the dependent variable that is predictable from the independent variables.
$R^{2}=1-\frac{\text{SS}_{\text{res}}}{\text{SS}_{\text{tot}}}$
where $\text{SS}_{\text{tot}}=\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$ measures the total variance in the observed data and $\text{SS}_{\text{res}}=\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ measures the variance not explained by the model.
3.4 Performance
We next provide simulation results using different dataset and models. On reporting and explaining the performance of each combination of model and dataset, we focus on the achievement of the three aforementioned purposes.
$E_{\text{coul}}$ computed using electrostatic features only on Dataset 1 and Dataset 2
We first investigate the effectiveness of electrostatic features in the prediction of the Coulombic energy $E_{\text{coul}}$. Note $E_{\text{coul}}$ is determined by the location and charges of the particles, which also determines the moments of the multipoles as the electrostatic features.
Table 1: $E_{\text{coul}}$ predicted using Electrostatic Features only on Dataset 1 and Dataset 2; D2-D1 indicates the difference of corresponding values of MSE and $R^{2}$ from Dataset 1 and Dataset 2.
$p$ $L$ $N_{f}(p,L)$ Dataset 1 Dataset 2 D2-D1
MSE MAPE $R^{2}$ MSE MAPE $R^{2}$ MSE MAPE $R^{2}$
1 0 4 0.641 0.328 0.351 0.682 0.532 0.331 0.041 0.204 -0.020
2 0 10 0.426 0.253 0.569 0.351 0.352 0.655 -0.074 0.100 0.086
3 0 20 0.352 0.206 0.643 0.252 0.318 0.753 -0.100 0.111 0.110
4 0 35 0.279 0.190 0.717 0.270 0.361 0.735 -0.009 0.172 0.018
0 1 9 0.699 0.354 0.292 0.663 0.548 0.349 -0.036 0.194 0.057
1 1 36 0.295 0.206 0.701 0.235 0.277 0.769 -0.060 0.070 0.068
2 1 90 0.284 0.175 0.712 0.154 0.224 0.848 -0.129 0.049 0.136
3 1 180 0.207 0.178 0.791 0.197 0.260 0.806 -0.009 0.082 0.016
4 1 315 0.203 0.157 0.794 0.224 0.320 0.780 0.021 0.163 -0.014
0 2 73 0.469 0.261 0.525 0.337 0.332 0.669 -0.131 0.071 0.144
1 2 292 0.249 0.182 0.748 0.144 0.201 0.859 -0.105 0.019 0.111
2 2 730 0.227 0.177 0.770 0.170 0.217 0.833 -0.057 0.040 0.063
3 2 1460 0.211 0.166 0.786 0.162 0.220 0.841 -0.049 0.054 0.055
4 2 2555 0.219 0.183 0.778 0.180 0.297 0.823 -0.039 0.114 0.045
0 3 585 0.465 0.265 0.529 0.287 0.283 0.719 -0.179 0.018 0.190
1 3 2340 0.513 0.220 0.480 0.255 0.259 0.750 -0.259 0.039 0.270
2 3 5850 0.469 0.230 0.525 0.215 0.240 0.789 -0.254 0.010 0.264
3 3 11700 0.450 0.220 0.544 0.269 0.248 0.736 -0.181 0.028 0.192
4 3 20475 0.451 0.221 0.543 0.241 0.323 0.764 -0.210 0.102 0.221
(a)                 (b)                 (c)
Figure 6: (a) $R^{2}$ of model model performance against number of features; (b) scatter plot for Dataset 1 ($p=4$, $L=1$ with total 315 features); (c) scatter plot for Dataset 2 ($p=1$, $L=2$ with total 292 features)
Table 1 reports the $E_{\text{coul}}$ prediction using electrostatic only on Dataset 1 and Dataset 2. The first important pattern we can observe from the table is that the MSE and $R^{2}$ obtained from Dataset 2 are essentially better than those obtained from Dataset 1 as the differences of MSE (D2-D1) are almost all negative and the difference of $R^{2}$ (D2-D1) is almost all positive. This pattern validates Purpose 2 indicating advantage of using larger dataset, particularly at capturing more overall variance in the target. We also noticed the difference of MAPE(D2-D1) is almost all positive, which apparently contradicts the results from MSE and $R^{2}$. In fact, this gives more information about the data distribution, i.e. the data in Dataset 2 but not in Dataset 1 might contain more smaller target values, yielding large percentage errors.
Also from Table 1, we can see on Dataset 1, MSE decreases and $R^{2}$ increase with the increment of $p$ and $L$, reaching optimized value at $p=4$ and $L=1$ with 315 features. Similarly on Dataset 2, MSE decreases and $R^{2}$ increases with the increment of $p$ and $L$ too, reaching optimized value at $p=1$ and $L=2$ with 292 features. This pattern validates the Purpose 1 showing the advantage of using more features. However, we should aware that optimized $p$ and $L$ are affected by the size of the training dataset. When the number of features are larger than number of samples, increasing number of features might bring negative impact such as overfitting.
We plot $R^{2}$ in Fig. 6(a) with horizontal axis as number of features $N_{f}$ from the third column of Table 1. The validation of Purpose 1 is shown since blue squares are lower than the red ones. The validation of Purpose 2 is also shown since markers of the same color contain an increasing pattern as number of features increase until there are too many features. We also plot the best cases on using both Datasets in Fig. 6(b,c) and we can see the distribution of the data points are essentially around the $\hat{y}=y$ line.
$E_{\text{coul}}$ computed using e-features and both features on Dataset 2
Table 2: $E_{\text{coul}}$ computed using e-features and both features on Dataset 2; the MSE and $R^{2}$ for using Topological Features only is 0.039 and 0.962 respectively.
$p$ $L$ $N(p,L)$ e-features both features difference
MSE $R^{2}$ MSE $R^{2}$ MSE $R^{2}$
1 0 4 0.682 0.331 0.039 0.961 -0.643 0.631
2 0 10 0.351 0.655 0.033 0.967 -0.318 0.312
3 0 20 0.252 0.753 0.044 0.956 -0.208 0.203
4 0 35 0.270 0.735 0.030 0.970 -0.240 0.235
0 1 9 0.663 0.349 0.030 0.970 -0.634 0.621
1 1 36 0.235 0.769 0.040 0.960 -0.195 0.191
2 1 90 0.154 0.848 0.044 0.956 -0.110 0.108
3 1 180 0.197 0.806 0.037 0.963 -0.160 0.156
4 1 315 0.224 0.780 0.045 0.955 -0.179 0.175
0 2 73 0.337 0.669 0.048 0.952 -0.289 0.283
1 2 292 0.144 0.859 0.062 0.939 -0.082 0.080
2 2 730 0.170 0.833 0.051 0.949 -0.119 0.116
3 2 1460 0.162 0.841 0.032 0.969 -0.131 0.128
4 2 2555 0.180 0.823 0.039 0.961 -0.141 0.138
0 3 585 0.287 0.719 0.047 0.953 -0.240 0.235
1 3 2340 0.255 0.750 0.050 0.951 -0.205 0.201
2 3 5850 0.215 0.789 0.042 0.958 -0.172 0.169
3 3 11700 0.269 0.736 0.032 0.968 -0.236 0.232
4 3 20475 0.241 0.764 0.037 0.964 -0.204 0.200
We next investigate the effectiveness of combined features. This investigation has two parts. The first part is using $E_{\text{coul}}$ as labels with Dataset 2 and the second part is using $E_{\text{solv}}$ as labels with Dateset 1 considering the availability of labels and the size of the datasets. The MSE and $R^{2}$ for different models are shown in Table 2. Note the MSE and $R^{2}$ for using topological features only is 0.039 and 0.962 respectively, which is already very effective. If we see the difference column, we can see that adding topological features fundamentally improves the model using electrostatic features only by decreasing MSE and increase $R^{2}$. This partially validates purpose 2. Note we did not report MAPE result here as it shows similar pattern as MSE does. If we consider topological features first, adding electrostatic features seems only occasionally improve the results. This is because the results from using topological features are too good for the effectiveness of electrostatic features to kicked in. Our next test case shows when the topological features are relatively less effective, adding the electrostatic features can significantly improve the performance. Before we move to that case, let’s linger on Fig. 7, which is the scatter plot of predicted values vs true values for using topological features, electrostatic feature and combined feature, for the observation that the improvement from (b) to (c) are very obvious while the improvement from (a) to (c), although not that significant, is still discernible.
(a)                 (b)                 (c)
Figure 7: scatter plots: (a) topological features only, (b) e-features only ($p=1$, $L=2$ with total 292 features), (c) both features ($p=4$, $L=0$ with total 35 features).
$E_{\text{solv}}$ computed using e-feature and both features on Dataset 1
As mentioned earlier, computing electrostatic solvation energies accurately is time consuming. Here we computed the solvation energies for the proteins from Dataset 1 using the MIBPB [20] solver. These solvation energies are used as labels for the neural network training and testing. A similar simulation using Dataset 1 are conducted and the results are reported in Table 3. Form the table, we can see that the improvement on using both features from using topological features only or from using electrostatic features are significant. If we focus on the difference column (both features - e-features only), we can see adding topological features completely improve the model using electrostatic features only by decreasing MSE and increase $R^{2}$. Meanwhile, the MSE and $R^{2}$ in the columns of using both features are obviously better than using topological feature only, which has MSE = 0.150 and $R^{2}$= 0.826. These results validate purpose 3, showing the advantage of using topological and electrostatic features combined. Again, we did not report MAPE result here as it shows similar pattern as MSE does.
Table 3: $E_{\text{solv}}$ computed using e-features and both features on Dataset 1; the MSE and $R^{2}$ for using Topological Features only is 0.150 and 0.826 respectively.
$p$ $L$ $N(p,L)$ e-features both features difference
MSE $R^{2}$ MSE $R^{2}$ MSE $R^{2}$
1 0 4 0.345 0.604 0.063 0.926 -0.282 0.323
2 0 10 0.271 0.690 0.090 0.895 -0.180 0.206
3 0 20 0.135 0.845 0.064 0.926 -0.071 0.081
4 0 35 0.121 0.861 0.065 0.925 -0.056 0.064
0 1 9 0.447 0.487 0.114 0.867 -0.333 0.380
1 1 36 0.148 0.830 0.077 0.911 -0.072 0.081
2 1 90 0.095 0.891 0.084 0.903 -0.012 0.012
3 1 180 0.111 0.873 0.105 0.878 -0.006 0.005
4 1 315 0.128 0.853 0.097 0.887 -0.031 0.034
0 2 73 0.214 0.754 0.092 0.892 -0.122 0.138
1 2 292 0.169 0.806 0.127 0.852 -0.041 0.045
2 2 730 0.164 0.812 0.128 0.851 -0.036 0.039
3 2 1460 0.212 0.757 0.138 0.839 -0.074 0.082
4 2 2555 0.218 0.750 0.165 0.808 -0.053 0.058
0 3 585 0.336 0.615 0.148 0.827 -0.188 0.213
1 3 2340 0.399 0.542 0.218 0.747 -0.181 0.205
2 3 5850 0.390 0.553 0.171 0.801 -0.219 0.249
3 3 11700 0.469 0.461 0.195 0.773 -0.275 0.312
4 3 20475 0.365 0.581 0.242 0.718 -0.123 0.137
The optimized case for using e-feature only happens at $p=1$ and $L=1$ with 90 features. Adding topological features improve the performance. We provide the scatter plot between reference values and predicted values for using topological features only, electrostatic features only, and both features in Fig. 8. The improvements of using two kind of features as shown in (c) can be seen as compared with topological features only in (a) and electrostatic features only in (b).
(a)                 (b)                 (c)
Figure 8: scatter plots: (a) topological features only, (b) e-features only ($p=2$, $L=1$ with total 90 features), (c) both features.
4 Software Dissemination
The code associated with this project is free and publically available on the GitHub deposit: https://github.com/atalha-cmd/Biophysics_ML.git maintained by SMU graduate student Md Abu Talha. It contains the Fortran code for generating the electrostatic features using Cartesian treecode, the Python code to generate topological features (a modified version as used in [19]), and the Python code for the neural network, which is originated from Elyssa Sliheet’s thesis work [66].
5 Conclusion
The project provides a novel approach to generate topological and electrostatic features combined to represent protein structures and charge distributions in a physical-informed multiscale and uniform way. Using these features, we can predict important protein properties such as the solvation energy and Coulombic energies. The numerical simulation validates our three purposes such that larger dataset, more electrostatic features, and combined features can improve the performance of the machine learning model. Other than predicting solvation energies and Coulomb energy, the resulting electrostatic and topological features potentially have broad usage as far as the proteins are needed to be represented in the forms of uniform and multiscale features for machine learning models. In the future, we will further investigate the role of reaction potential, which will be calculated using the much faster GB model as explained in the manuscript. We will also investigate Barycentric Lagrange treecode to generate electrostatic features as compared with the Cartesian treecode.

### v3

@@
 Abstract
-In this project, we provide a deep-learning neural network (DNN) based biophysics model to predict protein properties.
-The model uses multi-scale and uniform topological and electrostatic features generated using protein structural information and force field.
-The topological features are generated using the element specified persistent homology (ESPH) on a selection of heavy atoms or carbon atoms which contribute significantly in the protein biophysical properties.
-The electrostatic features are fast computed using a Cartesian treecode, modulated with treecode parameters for accuracy-cost tradeoff.
-These features are uniform in number for proteins with various sizes thus the broadly available protein structure database can be used to train the network.
-These features are also multi-scale thus the resolution and computational cost can be balanced by the users at the consideration of available computing powers.
-The machine learning simulations on two data sets with 4000+ and 17000+ protein structures showed the efficiency and fidelity of these features in representing the protein structure and force field for the predication of their biophysical properties such as electrostatic Coulombic energies and solvation energies.
-The feature generation algorithms have the potential as general tools in assisting machine learning based biophysical properties and function prediction for the broad biomolecules whose structures are obtained from both theoretical computing and experiments.
-keywords:
-Electrostatics; Poisson-Boltzmann; treecode; Deep Neural Network; Persistent Homology
+In this project, we present a deep neural network (DNN)-based biophysics model that uses multi-scale and uniform topological and electrostatic features to predict protein properties, such as Coulomb energies or solvation energies.
+The topological features are generated using element-specific persistent homology (ESPH) on a selection of heavy atoms or carbon atoms.
+The electrostatic features are generated using a novel Cartesian treecode, which adds underlying electrostatic interactions to further improve the model prediction.
+These features are uniform in number for proteins of varying sizes; therefore, the widely available protein structure databases can be used to train the network.
+These features are also multi-scale, allowing users to balance resolution and computational cost.
+The optimal model trained on more than 17,000 proteins for predicting Coulomb energy achieves MSE of approximately 0.024, MAPE of 0.073 and $R^{2}$ of 0.976.
+Meanwhile, the optimal model trained on more than 4,000 proteins for predicting solvation energy achieves MSE of approximately 0.064, MAPE of 0.081, and $R^{2}$ of 0.926, showing the efficiency and fidelity of these features in representing the protein structure and force field.
+The feature generation algorithms also have the potential to serve as general tools for assisting machine learning based prediction of protein properties and functions.
+[…]
+{tocentry}[Uncaptioned image]
 1 Introduction
 One of the overarching themes of biology is that structure determines function, which becomes more significant when we explore the biological world at the molecular level.
-With the advancement and availability of structure determination techniques such as X-ray crystallography [1, 2], NMR spectroscopy [3], cryo-electron microscopy [4], etc., there are rapidly growing 3D structures of protein, nucleic acids, and complex assemblies, etc.
-on publicly accessible repository such as Protein Data Bank (PDB) [5], GenBank [6], etc.
+With the advancement and availability of structure determination techniques such as X-ray crystallography Bragg:1912, Crowfoot:1935, NMR spectroscopy Aue:1976, cryo-electron microscopy Cheng:2015, etc., there are rapidly growing 3D structures of protein, nucleic acids, and complex assemblies, etc.
+on publicly accessible repository such as Protein Data Bank (PDB) Berman:2000, GenBank Benson:2012, etc.
 These structure determination techniques have strengths and limitations that make them suitable for different types research studies.
@@
 Cryo-electron microscopy which becomes more and more popular recently, excels in visualizing large, complex, and flexible biomolecular systems.
-The combination and integration of these methods due to their complementary natures help to reveal critical biomolecular functions embedded in the structures.
 In addition to experiments, computing approaches are also available for the prediction of protein structures and properties.
-The main stream of Machine Learning based protein prediction focuses on using amino acid sequence to predict protein structure or structure-related properties, resulting in the trending protein language models (PLM) [7, 8], which is a transformer-based model inspired by Natural Language Processing (NLP) with wide range of applications such as protein pKa prediction [9], protein structure prediction [10], protein design [11], protein evolution [12], etc.
-Recently, the appearance of AlphaFold developed by Google DeepMind revolutionarily reshaped the structures determination techniques.
-Based on known structures, AlphaFold learned to predict protein’s structures from sequence with very high accuracy as shown by their victory by a large margin in Critical Assessment of Techniques for Protein Structure Prediction (CASP 14-16), AlphaFold Protein Structures Database now has over 200 million protein structure predictions [13].
-Both PLM and AlphaFold take protein sequences as inputs to predict structures or embeddings which keep important features, without touching the prediction of protein functions from protein structures.
-The machine learning methods potentially can revolutionize the traditional structure-based molecular modeling, which uses brownian dynamics, molecular dynamics, Monte Carlo simulations, etc.
-to predict protein properties and functions from protein structures.
+The main stream of Machine Learning (ML) based protein prediction focuses on using amino acid sequence to predict protein structure or structure-related properties, resulting in the trending protein language models (PLM) Bepler:2021, Weissenow:2025, which is a transformer-based model inspired by Natural Language Processing (NLP).
+PLM has a wide range of applications such as protein pKa prediction Xu:2025, protein structure prediction Lin:2023, protein design Madani:2023, protein evolution Zhang:2024, etc.
+Recently, the appearance of AlphaFold developed by Google DeepMind revolutionarily reshaped the structures determination techniques Abramson:2024.
+Both PLM and AlphaFold take protein sequences as inputs to predict structures or embeddings which keep important features.
 However, the development of ML methods in protein property predictions is hindered by the difficulty to represent protein structures and force fields using reliable features.
-To this end, the structure data, which is largely variant from protein to protein, needs to be converted to uniform size of features in representing structures to be loaded into the machine learning models such as regression models, random forest, deep neural network etc.
-Successful cases are the use of graph convolutional network (GCN) for predicting protein functions by features extracted from a protein language model and protein structure [14], protein-DNA binding specificity prediction using a wide range of features from geometry [15, 16], protein-ligand binding affinity prediction using topological features, [17, 18].
-In this paper, we provide a novel way to generate electrostatic features.
-Combining electrostatic features with the topological features generated using a modification of the persistent homology approach used in [19], we represent the protein structures and associated force field in a uniform and multiscale fashion.
-With these features, we train a deep-learning neural network (DNN) to predict protein Coulombic energy and electrostatic solvation energy efficiently and accurately.
-Here the Solvation energies as labels are generated by solving the Poisson-Boltzmann (PB) model using a second-order Matched Interface and Boundary (MIB) method based PB solver [20].
-There are recent research using Physics Informed Neural Network (PINN) to solve the Poisson-Boltzmann model [21, 22] while the present work is rather different as a Data-driven method with uniform and multi-scale features.
-In addition to these particular application in predicting Coulombic energy and electrostatic solvation energy, the feature generation algorithm we have developed potentially has wider application for other machine learning models as far as the protein structures and force field are required to be represented uniformly for proteins with various sizes.
-This manuscript is constructed in the following way.
-After this section of introduction is the section of models and algorithms, followed by the section of results of simulation.
-The manuscript is ended with a conclusion section.
-2 Models and Algorithms
-We introduce the models and algorithms involved in this work.
-For protein’s structure-property relation we focus on the atomic description and the electrostatic interactions.
-The protein properties involved in this paper are Coulombic energy $E_{\text{coul}}$ and electrostatic solvation energy $E_{\text{solv}}$.
-For a protein, the Coulombic energy $E_{\text{coul}}$ is the energy it takes to displace atomic charges of the protein from infinity to their present positions (i.e.
-atomic centers from the PDB file) while the electrostatic solvation energy $E_{\text{solv}}$ is the energy it takes to solvate the protein into the solvent (e.g.
-water).
-Formal quantitative definition of these two protein properties will be given with mathematical models.
-For the protein-solvent interaction, we use the implicit solvent model in which water is represented as a continuum for the reduced computational cost.
-In a comparison perspective, we introduce the two most popular implicit solvation models, namely the Poisson-Boltzmann model and the Generalized Born’s model, which are used to generate the labels or features.
-Following that, we introduce the persistent homology and how we use it to produce the topological features.
-Finally we introduce how to use the Cartesian treecode, whose modification evolves into the algorithm to generate the electrostatic features.
+To this end, the structure data, which is largely variant from protein to protein, needs to be converted to uniform size of features to be loaded into the ML models.
+Successful cases are the use of graph convolutional network (GCN) for predicting protein functions by features extracted from a protein language model and protein structure Gligorijevic:2021, protein-DNA binding specificity prediction using a wide range of features from geometry Sagendorf:2024, Mitra:2025, protein-ligand binding affinity prediction using topological features, Wang:2017, Wee:2022, molecular structure representation using persistent diagram Townsend:2020, etc.
+However, the important electrostatic interactions are often ignored in the formation of the protein features because of the challenges from long-range and pairwise nature of the electrostatic interactions.
+In this paper, we provide a novel approach that utilizes the far-field expansion in the Cartesian treecode algorithm Li:2009 to generate electrostatic features.
+To this end, the charges at the atomic centers are represented by multipole moments at the cluster centers, which inherit the key idea of the treecode algorithm Li:2009: replacing pairwise particle-particle interactions with particle-cluster interaction.
+By combining electrostatic features with the topological features generated using a modification of the persistent homology approach used by Cang and Wei Cang:2017, we represent the protein structures and the associated force field in a uniform and multiscale fashion.
+With these features, we train a deep neural network (DNN) to predict protein Coulombic energy and electrostatic solvation energy efficiently and accurately.
+Here, the Coulomb Energy and Solvation energy used as labels are generated by pair-wise Coulombic interactions and the surrogate model for solving the Poisson-Boltzmann (PB) model using a second-order Matched Interface and Boundary (MIB) method based PB solver Chen:2011.
+Apart from the recent research using Physics Informed Neural Network (PINN) to solve the Poisson-Boltzmann model Achondo:2025, Chen:2025, the present work is fundamentally different as a data-driven method with uniform and multiscale features.
+There exist machine learning methods for the prediction of solvation energy under the framework of the PB model.
+For examples, Huang et al.
+developed a neural-network formulation for solving the dielectric-boundary PB equation by minimizing a variational electrostatic free-energy functional, enabling direct prediction of PB electrostatic solvation energies while preserving the physical structure of the continuum model Huang:2026; Röcken et al.
+proposed the ReSolv framework trained using experimental hydration free energies and ab initio data, achieving near experimental accuracy while accelerating solvation energy prediction Rocken:2024; Ferraz-Caetano et al.
+introduced an explainable supervised learning approach based on ensemble regression models that predicts solvation energies from physicochemical descriptors FerrazCaetano:2024; Wu and Luo presented SPACIER Wu:2025, a large-scale benchmark dataset of solvation energies grounded computed by using AMBER-PBSA Case2023AmberTools.
+The major advantage of the proposed method is at the feature generation independent of the PB framework, thus in addition to this particular applications in predicting Coulombic energy and PB-based electrostatic solvation energy, the feature generation algorithm developed in this work potentially has broader applications to many other ML models, as far as protein structures and force fields are the inputs represented uniformly across proteins of varying sizes.
+This manuscript is organized as follows.
+Following this introduction, we present the methods section, which includes both the modeling framework and numerical algorithms, followed by the simulation results.
+The manuscript ends with a section of conclusion.
+2 Methods
+The objective of this project is to design uniform and multiscale features that represent protein structure and electrostatics, and then to use ML models, such as DNN, to evaluate the effectiveness of these features.
+The models are trained using labels such as solvation energy computed from surrogate models, e.g.
+the numerically solved Poisson-Boltzmann model.
+We hereby introduce the models and algorithms involved in this work.
+To study protein structure–property relationships, we focus on the atomic description and the electrostatic interactions.
+The protein properties considered in this paper are Coulomb energy $E_{\text{coul}}$ and electrostatic solvation energy $E_{\text{solv}}$.
+For a protein, the Coulomb energy $E_{\text{coul}}$ is the energy required to move atomic charges from infinity to their present positions(i.e.
+atomic centers specified in the Protein Data Bank (PDB) file.
+The electrostatic solvation energy $E_{\text{solv}}$ is the electrostatic free energy required to solvate the protein in a solvent, such as water.
+Formal quantitative definitions of these two properties are provided through mathematical models.
+For the protein-solvent interaction, we use the implicit solvent model in which water is treated as a continuum in order to reduce computational cost.
+From a comparative perspective, we introduce the two most popular implicit solvation models, namely the Poisson-Boltzmann model and the Generalized Born’s model, which are used to generate labels or features.
+We then introduce persistent homology and describe how it is used to construct topological features.
+Finally, we present the Cartesian treecode Li:2009 algorithm, and its modification into a point-multipole representation, which forms the basis for generating electrostatic features.
+In this project, we adopt a combination of electrostatic features and topological features for the following reasons.
+Electrostatic features provide a uniform and multiscale representation of the atomic charges, accounting for both charge quantities and their locations.
+Topological features, on the other hand, encode information derived from the atomic centers of selected atoms, such as alpha-carbon atoms or heavy atoms.
+These features reveal the intrinsic topological invariants, which are not directly represented by electrostatic features.
+Together, these two types of features efficiently represent charge distributions, their locations, and underlying topological information.
 2.1 The Poisson-Boltzmann model
@@
 1(a) a protein is represented by a collection of $N_{c}$ spherical atoms with centered partial charges.
-The molecular surface $\Gamma$ (also known as the solvent excluded surface) is defined by the trace of a water molecule represented by a red sphere rolling on contacting with the protein atoms.
+The molecular surface $\Gamma$ (also known as the solvent excluded surface Richards:1977, Connolly:1985) is defined by the trace of a water molecule represented by a red sphere rolling on contacting with the protein atoms.
 The Poisson-Boltzmann model is shown in Fig.
-1(b), where the molecular surface $\Gamma$ divides the entire computational domain $\Omega$ into the protein domain $\Omega_{1}$ with dielectric constant $\varepsilon_{1}$ and atomic charges $q_{k}$ located at ${\bf x}_{k},k=1\,{:}\,N_{c}$, and the solvent domain $\Omega_{2}$ with dielectric constant $\varepsilon_{2}$ and dissolved salt ions  [23, 24].
+1(b), where the molecular surface $\Gamma$ divides the entire computational domain $\Omega$ into the protein domain $\Omega_{1}$ with dielectric constant $\varepsilon_{1}$ and atomic charges $q_{k}$ located at ${\bf x}_{k},k=1\,{:}\,N_{c}$, and the solvent domain $\Omega_{2}$ with dielectric constant $\varepsilon_{2}$ and dissolved mobile ions.
 Assuming a Boltzmann distribution for the ion concentration, and considering the case of two ion species with equal and opposite charges (e.g.
-Na+, Cl-), in the limit of weak electrostatic potential one obtains the linearized Poisson-Boltzmann (PB) model [25],
-$\varepsilon_{1}\nabla^{2}\phi_{1}(\textbf{x})$ $=-\sum_{k=1}^{N_{c}}q_{k}\delta(\textbf{x}-\textbf{x}_{k}),\quad{\bf x}\in\Omega_{1},$ (1a)
-$\varepsilon_{2}\nabla^{2}\phi_{2}(\textbf{x})$ $=\kappa^{2}\phi_{2}(\textbf{x}),\quad{\bf x}\in\Omega_{2},$ (1b)
+Na+, Cl-), in the limit of weak electrostatic potential one obtains the linearized Poisson-Boltzmann (PB) model Baker:2005,
+$\varepsilon_{1}{\rm NewA}^{2}\phi_{1}(\textbf{x})$ $=-{4\pi}\sum_{k=1}^{N_{c}}q_{k}\delta(\textbf{x}-\textbf{x}_{k}),\quad{\bf x}\in\Omega_{1},$ (1a)
+$\varepsilon_{2}{\rm NewA}^{2}\phi_{2}(\textbf{x})$ $=\kappa^{2}\phi_{2}(\textbf{x}),\quad{\bf x}\in\Omega_{2},$ (1b)
 $\phi_{1}(\textbf{x})$ $=\phi_{2}(\textbf{x}),\quad\varepsilon_{1}\frac{\partial\phi_{1}}{\partial n}({\bf x})=\varepsilon_{2}\frac{\partial\phi_{2}}{\partial n}({\bf x}),\quad{\bf x}\in\Gamma,$ (1c)
@@
 Figure 1: Implicit solvent models.
-(a) The Molecular Surface $\Gamma$: the trace of the solvent probe (shown in red) when it is rolled on contacting the spherical atoms of the protein (shown in green with a centered charge); (b) The Poisson–Boltzmann Model: two domains $\Omega_{1}$ (dielectric constant $\epsilon_{1}$ with partial charges as weighted summation of delta functions) and $\Omega_{2}$ (dielectric constant $\epsilon_{2}$ with mobile ions modeled by Boltzmann distribution) are separated by the molecular surface $\Gamma$; (3) The Generalized Born Model: the protein is represented as a collection of $N_{c}$ spherical atoms centered at ${\bf x}_{k}$ with charge $q_{k}$ and Born Radius $R_{k}$ (only the $k$th atom is shown).
+(a) The Molecular Surface $\Gamma$: the trace of the solvent probe (shown in red) when it is rolled on contacting the spherical atoms of the protein (shown in green with a centered charge); (b) The Poisson–Boltzmann Model: two domains $\Omega_{1}$ (dielectric constant $\epsilon_{1}$ with partial charges as weighted summation of delta functions) and $\Omega_{2}$ (dielectric constant $\epsilon_{2}$ with mobile ions modeled by the Boltzmann distribution) are separated by the molecular surface $\Gamma$; (3) The Generalized Born Model: the protein is represented as a collection of $N_{c}$ spherical atoms centered at ${\bf x}_{k}$ with charge $q_{k}$ and Born Radius $R_{k}$ (only the $k$th atom is shown).
 The PB model governs the electrostatic potential $\phi$ in the entire space.
@@
 However, there are many challenging issues on properly obtaining the field and energy (e.g.
-definition of field on molecular surface $\Gamma$ [26]).
+definition of field on molecular surface $\Gamma$ Gilson:1993, Geng:2011).
 Our attention for this project is on the energy side as described below.
@@
 $E_{\text{free}}=\frac{1}{2}\int_{\Omega}\rho({\bf x})\phi({\bf x})d{\bf x}=\frac{1}{2}\sum\limits_{k=1}^{N_{c}}q_{k}\phi({\bf x}_{k})=\frac{1}{2}\sum\limits_{k=1}^{N_{c}}q_{k}(\phi_{\text{reac}}({\bf x}_{k})+\phi_{\text{coul}}({\bf x}_{k}))=E_{\text{solv}}+E_{\text{coul}}$ (2)
-where $\rho({\bf x})=\sum\limits_{k=1}^{N_{c}}q_{k}\delta(\textbf{x}-\textbf{x}_{k})$ is the charge density as a sum of partial charges weighted delta function and the $E_{\text{solv}}=\frac{1}{2}\sum\limits_{k=1}^{N_{c}}q_{k}\phi_{\text{reac}}({\bf x}_{k})$ term is the electrostatic solvation energy.
-The $\phi_{\text{reac}}({\bf x}_{k})$ is the reaction potential at ${\bf x}_{k}$ as the remaining component when Coulomb potential $\phi_{\text{coul}}({\bf x}_{k})=\sum\limits_{j=1,j\neq k}^{N_{c}}\frac{q_{j}}{\epsilon_{1}|{\bf x}_{k}-{\bf x}_{j}|}$ is taken away from the total electrostatic potential $\phi({\bf x}_{k})$.
+where $\rho({\bf x})=\sum\limits_{k=1}^{N_{c}}q_{k}\delta(\textbf{x}-\textbf{x}_{k})$ is the charge density as a sum of partial charges weighted delta functions and the $E_{\text{solv}}=\frac{1}{2}\sum\limits_{k=1}^{N_{c}}q_{k}\phi_{\text{reac}}({\bf x}_{k})$ term is the electrostatic solvation energy.
+The $\phi_{\text{reac}}({\bf x}_{k})$ is the reaction potential at ${\bf x}_{k}$ as the remaining component when Coulomb potential $\phi_{\text{coul}}({\bf x}_{k})=\sum\limits_{j=1,j\neq k}^{N_{c}}\frac{q_{j}}{\epsilon_{1}|{\bf x}_{k}-{\bf x}_{j}|}$ is taken away from the electrostatic potential $\phi({\bf x}_{k})$.
 Solving the PB model numerically by grid-based methods is challenging because (1) the protein is represented by singular point charges, (2) the molecular surface is geometrically complex, (3) the dielectric constant is discontinuous across the surface, (4) the domain is unbounded.
-To overcome these numerical difficulties, several finite difference interface methods have been developed.
-By assuming that the interface is aligned with a mesh line, a jump condition capture scheme has been developed in [27].
-Based on a Cartesian grid, the immersed interface method (IIM) [28] has been applied to solve the PB equation [29, 30], in which the jump conditions can be rigorously enforced based on Taylor expansions.
-For the purpose of dealing with arbitrarily shaped dielectric interfaces based on a simple Cartesian grid, a matched interface and boundary (MIB) PB solver [31, 32, 33, 20, 34] has been developed through rigorous treatments of geometrical and charge singularities.
-Alternatively, boundary element methods (BEM) for the PB model were developed later [35, 36, 37, 38, 39, 40, 41, 42] with several inherent advantages, (1) only the molecular surface is discretized rather than the entire solute/solvent volume, (2) the atomic charges are treated analytically, (3) the interface conditions are accurately enforced, (4) the far-field boundary condition is imposed analytically, (5) the matrix-vector product in iterative methods can be accelerated by fast algorithms [43, 44, 45, 46].
-In this paper, we computed the solvation energies using the MIBPB solver [20] since accuracy is of critical importance for the generation of labels ($E_{\text{solv}}$) that will be used to train the neural network.
+To overcome these numerical difficulties, many finite difference interface methods have been developed Chern:2003, LeVeque:1994, Qiao:2006.
+For the purpose of dealing with arbitrarily shaped dielectric interfaces based on a simple Cartesian grid, a matched interface and boundary (MIB) PB solver Zhou:2006a, Yu:2007, Geng:2007, Chen:2011, Geng:2017 has been developed through rigorous treatments of geometrical and charge singularities.
+Alternatively, boundary element methods (BEM) for the PB model were developed later Zauhar:1988, Juffer:1991, Liang:1997, Boschitsch:2002, Bordner:2003, Lu:2007, Geng:2013b, Chen:2023 with several inherent advantages, (1) only the molecular surface is discretized rather than the entire solute/solvent volume, (2) the atomic charges are treated analytically, (3) the interface conditions are accurately enforced, (4) the far-field boundary condition is imposed analytically, (5) the matrix-vector product in iterative methods can be accelerated by fast algorithms Barnes:1986, Greengard:1987, Li:2009, Tausch:2003.
+In this paper, we computed the solvation energies using the MIBPB solver Chen:2011 since accuracy is of critical importance for the generation of labels ($E_{\text{solv}}$) that will be used to train the neural network.
 2.2 Generalized Born’s Model
@@
 Although we did not use the GB model to produce any numerical results for this project, we describe the GB model for its relationship with the PB model and its potential usage in electrostatic interactions.
-In fact, the GB model can calculate the electrostatic solvation energy which can be used either as the label or as a feature [47, 48] in our machine learning framework.
+In fact, the GB model can calculate the electrostatic solvation energy which can be used either as the label or as a feature Chen:2021, Chen:2024 in our ML framework.
 It can also produce the reaction potentials at the atomic center, which can be used to generate the second level of electrostatic features as shown in Fig.
-3, which will be explained within that content.
-To derive the GB model, we first adopt some classic results for electrostatics from Jackson’s textbook [49] (sec.
-1.11).
-Using the analogue from discrete point charge distribution for continuous charge density distribution, the electrostatic free energy of a charged system with density distribution $\rho({\bf r})$ takes the form
-$E_{\text{free}}=\frac{1}{2}\int\rho({\bf r})\phi({\bf r})d{\bf r}=\frac{1}{8\pi}\int{\bf E}({\bf r})\cdot{\bf D}({\bf r})d{\bf r}$
-by using the Poisson equation $\nabla^{2}\phi=\rho({\bf r})/\epsilon$, the electric field ${\bf E}=-\nabla\phi$, and electric displacement ${\bf D}=\epsilon{\bf E}$.
-Now consider assembling a charge $q_{i}$ to the center of a sphere of radius $a_{i}$ and the center is located at the origin.
-The sphere separates the domain with $\epsilon_{1}$ for $r<a_{i}$ and $\epsilon_{2}$ for $r>a_{i}$.
-The assembly takes the energy
-$E_{\text{free,}i}=\frac{1}{8\pi}\int\frac{{\bf D}\cdot{\bf D}}{\epsilon}d{\bf r}\approx\frac{1}{8\pi}\int_{r<a_{i}}\frac{q_{i}^{2}}{r^{4}\epsilon_{1}}d{\bf r}+\frac{1}{8\pi}\int_{r>a_{i}}\frac{q_{i}^{2}}{r^{4}\epsilon_{2}}d{\bf r}$ (3)
-where the Coulomb field approximation ${\bf D}=\displaystyle\frac{q_{i}{\bf r}}{r^{3}}$ is used.
-From a slightly different perspective as in Eq.
-(2), the electrostatic solvation energy is the energy difference when $\epsilon$ for $r>a_{i}$ changes from $\epsilon_{1}$ (unsolvated) to $\epsilon_{2}$ (solvated) given as
-$E_{\text{solv},i}=\frac{1}{8\pi}(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}})\int_{r>a_{i}}\frac{q_{i}^{2}}{r^{4}}d{\bf r}=(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}})\frac{q_{i}^{2}}{2a_{i}}.$ (4)
-This result is consistent with the Poisson-Boltzmann model of a solvated spherical cavity with centered charges as summarized in [50], which is a special case from Kirkwood’s derivation of a series of spherical harmonics for a spherical cavity containing arbitrary multiple charges [51].
-If we treat the molecule as a collection of spherical atoms, using Eq.
-(4) the total electrostatic solvation energy is
+4 to be explained within that content.
+Under the GB model, the total electrostatic solvation energy is
 $E_{\text{solv}}=\frac{1}{2}\left(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}}\right)\left(\sum_{i=1}^{N}\frac{q_{i}^{2}}{a_{i}}+\sum_{j\neq i}^{N}\frac{q_{i}q_{j}}{r_{ij}}\right)\approx\frac{1}{2}(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}})\sum_{i=1}^{N}\sum_{j=1}^{N}\frac{q_{i}q_{j}}{f_{ij}^{\text{GB}}}$ (3)
-where $r_{ij}$ is the distance between the atomic centers of atoms $i$ and $j$.
-Here the first equation of $E_{\text{solv}}$ has two terms: a sum of individual Born’s terms and pairwise Coulombic terms.
-Note the actual molecule is better represented by a dielectric interface (e.g.
-solvent excluded surface (SES)) which separates inside domain $\Omega_{1}$ and outside domain $\Omega_{2}$.
-The charge $q_{i}$ is located at the center ${\bf r}_{i}$ of a sphere with radius $a_{i}$.
-We assume $\Omega_{1}$ contains all these spheres.
-The model using dielectric interface requires a step further as been approximated in the second equation in Eq.
-(5).
-Here $f_{ij}^{\text{GB}}$ is the effective Born radii ($i=j$) or effective interaction distance ($i\neq j$).
-Assuming the effective Born Radii $R_{i}$’s are obtained, a popular estimation of $f_{ij}^{\text{GB}}$ is:
-$f_{ij}^{\text{GB}}(r_{ij})=(r_{ij}^{2}+R_{i}R_{j}e^{\frac{-r_{ij}^{2}}{4R_{i}R_{j}}})^{\frac{1}{2}}$
-Note $R_{i}$ depends not only on $a_{i}$, but also on radii and relative positions of all other atoms.
-The estimation of effective Born radii is an active research area [52].
-From Poisson-Boltzmann theory, the perfect Born radii is given as
-$R_{i}=\frac{1}{2}(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}})\frac{q_{i}}{E^{\text{PB}}_{\text{solv},i}}$
-with $E^{\text{PB}}_{\text{solv},i}$ as the electrostatic solvation energy from the PB model with all except $i$th charge muted.
+where $r_{ij}$ is the distance between the atomic centers of atoms $i$ and $j$, $f_{ij}^{\text{GB}}$ is the effective Born radii ($i=j$) or effective interaction distance ($i\neq j$).
 This PB-GB relationship can be seen in Fig.
 1(c), in which the electrostatic solvation energy of the sphere equals that of the molecule, both with a charge at the atomic center.
-There are many approaches to approximate $R_{i}$ and a typical one is the volume integration via quadrature as
-$R_{i}^{-1}=\frac{1}{4\pi}\int_{\Omega_{2}}\frac{q_{i}^{2}}{r^{4}}d{\bf r}=\frac{q_{i}^{2}}{a_{i}}-\frac{1}{4\pi}\int_{\Omega_{1}/B({\bf r}_{i},a_{i})}\frac{q_{i}^{2}}{r^{4}}d{\bf r}.$
+Some details about the GB model can be found in the supplement material.
 2.3 Topological Features
@@
 Formally, A $k$-simplex is the convex hull of $k+1$ affinely independent points $v_{0},v_{1},\dots,v_{k}$:
-$\sigma^{k}=\left\{\sum_{i=0}^{k}\lambda_{i}v_{i}\;\middle|\;\lambda_{i}\geq 0,\sum\lambda_{i}=1\right\}$
+$\sigma^{k}=\left\{\sum_{i=0}^{k}\lambda_{i}v_{i}\;\middle|\;\lambda_{i}\geq 0,\sum\lambda_{i}=1\right\}$ (4)
 Each subset of $m+1$ of these vertices forms an $m$-face of the simplex, which is itself an $m$-simplex.
 The boundary of a $k$-simplex $v_{0},\dots,v_{k}]$ is:
-$\partial_{k}[v_{0},v_{1},\dots,v_{k}]=\sum_{i=0}^{k}(-1)^{i}[v_{0},\dots,\hat{v}_{i},\dots,v_{k}]$
+$\partial_{k}[v_{0},v_{1},\dots,v_{k}]=\sum_{i=0}^{k}(-1)^{i}[v_{0},\dots,\hat{v}_{i},\dots,v_{k}]$ (5)
 where $\partial_{k}$ is the boundary operator from $k$-chains to (k1)(k-1)-chains and $\hat{v}_{i}$ means vertex $v_{i}$ is omitted.
 Here a $k$-chain is a formal linear combination of k-dimensional simplexes in a simplicial complex.
-With the concepts of simplex, face, boundary, chain, etc., we can formally define that a simplicial complex $K$ is a collection of simplices such that:
+With the concepts of simplex, face, boundary, chain, etc., we define that a simplicial complex $K$ is a collection of simplices such that:
   1. 1.
@@
     The intersection of any two simplices is either empty or a common face.
-Thus, each simplicial complex corresponds to a topological space by “filling in” the simplexes and gluing them together along shared faces.
+Thus, each simplicial complex corresponds to a topological space by filling in the simplexes and gluing them together along shared faces.
+Let $C_{k}(K)$ be the group of $k$-chains, and let
+$\cdots\xrightarrow{\partial_{k+1}}C_{k}(K)\xrightarrow{\partial_{k}}C_{k-1}(K)\xrightarrow{\partial_{k-1}}\cdots$ (6)
+be the sequence of boundary maps $\partial_{k}$, the $k$-homology group of $K$ is defined as:
+$H_{k}(K)=\frac{\ker(\partial_{k})}{\operatorname{im}(\partial_{k+1})}$ (7)
+where
+  • $\ker(\partial_{k})$ is the group of $k$-cycles: $Z_{k}=\{c\in C_{k}\mid\partial_{k}(c)=0\}$
+  • $\operatorname{im}(\partial_{k+1})$ is the group of $k$-boundaries: $B_{k}=\{\partial_{k+1}(c)\mid c\in C_{k+1}\}$
+Hence, the $k$-th homology group is the group of cycles modulo boundaries:
+$H_{k}=Z_{k}/B_{k}.$ (8)
+Taking $H_{1}$ as an example, $Z_{1}=\ker(\partial_{1})$ collects 1-chains (formal sum of edges) without boundary points thus general loops, while $B_{1}=\operatorname{im}(\partial_{2})$ collects boundaries of 2-chains (a collection of triangles) that bound a filled-in region thus trivial holes, $H_{1}=Z_{1}/B_{1}$ is then the collection of actual holes in the space.
 2.3.2 Persistent Homology
 However, the simplicial homology of a complex associated to a point set at a particular proximity parameter $\epsilon$ is insufficient to characterize the signal or noise.
-This calls for the ideas of filtration and persistent homology, which can identify and connect complexes at different level of complexity and record the appearance and disappearance of the homology group.
-The importance of persistence is that given a parameterized family of spaces, topological features which persist over a significant parameter ranges are considered as signal while those with short-lived features are treated as noise [53].
+This calls for the ideas of filtration and persistent homology edelsbrunner:2000, zomorodian:2005, which can identify and connect complexes at different level of complexity and record the appearance and disappearance of the homology group.
 To this end, we introduce the concept of filtration, which is a sequence of nested simplicial complexes:
-$\emptyset=K_{0}\subseteq K_{1}\subseteq\cdots\subseteq K_{n}=K.$
+$\emptyset=K_{0}\subseteq K_{1}\subseteq\cdots\subseteq K_{n}=K.$ (9)
 Each $K_{i}$ is a simplicial complex, and the sequence grows over time or scale.
-Instead of computing just the $k$-homology group $H_{k}(K)$, which is an algebraic object capturing the $k$-dimensional holes in a topological space represented by a simplicial complex $K$, we compute $H_{k}(K_{i})$ for each $i$ and track features as the birth, death, and persistence of a homology class.
-Let $C_{k}(K)$ be the group of $k$-chains, and let
-$\cdots\xrightarrow{\partial_{k+1}}C_{k}(K)\xrightarrow{\partial_{k}}C_{k-1}(K)\xrightarrow{\partial_{k-1}}\cdots$
-be the sequence of boundary maps, the $k$-homology group of $K$ is defined as:
-$H_{k}(K)=\frac{\ker(\partial_{k})}{\operatorname{im}(\partial_{k+1})}$
-where
-  • 1.
-    $\ker(\partial_{k})$ is the group of $k$-cycles: $Z_{k}=\{c\in C_{k}\mid\partial_{k}(c)=0\}$
-  • 2.
-    $\operatorname{im}(\partial_{k+1})$ is the group of $k$-boundaries: $B_{k}=\{\partial_{k+1}(c)\mid c\in C_{k+1}\}$
-Hence, the $k$-th homology group is the group of cycles modulo boundaries:
-$H_{k}=Z_{k}/B_{k}$
-Taking $H_{1}$ as an example, $Z_{1}=\ker(\partial_{1})$ collects 1-chains (formal sum of edges) without boundary points thus general loops, while $B_{1}=\operatorname{im}(\partial_{2})$ collects boundaries of 2-chains (a collection of triangles) that bound a filled-in region thus trivial holes, $H_{1}=Z_{1}/B_{1}$ is then the collection of actual holes in the space.
+Instead of computing just the $k$-homology group $H_{k}(K)$, which is an algebraic object capturing the $k$-dimensional holes in a topological space represented by a simplicial complex $K$, we compute $H_{k}(K_{i})$ for each $i$, the index for time or scale parameters.
+As $i$ or its corresponding scale parameter grows, the duration, as measured by the change of $i$ or its corresponding scale parameter, of the homology class $H_{k}(K_{i})$ between its appearance and disappearance is called persistence.
+The importance of persistence is that given a parameterized family of spaces, topological features which persist over a significant parameter ranges are considered as signal while those with short-lived features are treated as noise Ghrist:2007.
+In practice, the persistent homology is computed using the pipeline that involves steps of turning data into a sequence of spaces (a filtration), tracking the birth and death of topological features using algebra, implementing using linear algebra tools, and visualization.
+The pipeline is realized by many established software, e.g.
+GUDHI Boissonnat:2014, as our choice.
+We next introduce the barcode, persistent diagram, and persistent image as approaches to visualize the persistent homology.
+These approaches are different in format and visualization but are equivalent mathematically.
+Following the definition, we provide the algorithms to extract topological features from barcode and persistent image.
+2.3.3 Barcode, Persistence Diagram and Persistence Image
+Persistent homology captures the evolution of topological features throughout a filtration process, which could be encoded in the forms of barcode, persistent diagram, and persistent image.
+A barcode for dimension $k$ is the multiset $\{[b_{j},d_{j})\}_{j}$, where each interval corresponds to a $k$-dimensional homology class with index $j$, which appears at $b_{j}$ and disappears at $d_{j}$.
+In the picture of a barcode, each class is visualized as a bar starting at $b_{j}$ and ending at $d_{j}$.
+Thus the barcode is a graphical representation of persistent homology as a collection of horizontal line segments in a plane whose horizontal axis corresponds to the filtrition parameter and whose vertical axis represents an (arbitrary) ordering of homology generators such as $H_{0}$, $H_{1}$, $H_{2}$, etc.
+Different colors can be used to distinguish different dimensions.
+A persistent diagram represents each homology class as a point $(b_{j},d_{j})$ in the birth–death plane, where $b_{j}$ and $d_{j}$ denote the filtration parameters at which the class appears and disappears.
+The persistent diagram thus is a scatter plot of points, where distance from the diagonal measures their significance.
+From their definitions, we can tell barcode and persistence diagram carry equal amount of information.
+We therefore only use barcode to generate topological features.
+The persistent image, which is built on top of persistent diagram to generate fixed-size 2-d image, is readily used as features or inputs for the convolution neural network.
+Details about persistent image and its formal mathematical definition can be found in supplementary materials.
 2.3.4 Feature generation from Barcode
-To make the topological features uniform and physically informed, we choose to use Element Specific Persistent Homology (ESPH) to extract topological features at different level of complexity.
-ESPH has been used for successfully predicting protein-ligand binding [19].
-The main idea is to select atom elements, which form the cloud of points, using target problem related physical, chemical, and biological information.
-In this project, since solvation energy and Coulomb energy are highly related to the amino acid chains, we use two collections of cloud of points.
-One is the collection of all Carbon atoms, which form the backbone of amino acids.
-The other is the collection of all heavy atoms {C, N, O, S}, whose locations and interconnections are also critical in protein structures.
-Once the collections of cloud of points are determined and selected, the barcode [53] carrying persisting homology information of the cloud of points are produced.
-Here the barcode is a graphical representation of persistent homology as a collection of horizontal line segments in a plane whose horizontal axis corresponds to the proximity parameter and whose vertical axis represents an (arbitrary) ordering of homology generators such as $H_{0}$, $H_{1}$, $H_{2}$, etc.
-We use the software GUDHI [54] to implement this task.
-The produced barcode is then used by the algorithm below to generate the vector of topological features.
-Figure 2: The barcode generated by the collection of cloud points using heavy atoms in proteins (a) 1a1e and (b) 1pb9.
-Figure 2 gives the barcode for protein 1a1e, the C-SRC Tyrosine Kinase, and protein 1pb9, the N-methyl-D aspartate receptor, using heavy atoms.
+To make the topological features uniform and physically informed, we choose to use Element-Specific Persistent Homology (ESPH) to extract topological features at different levels of complexity.
+ESPH has been successfully applied to the prediction of protein–ligand binding.
+Cang:2017.
+The core idea is to select atomic elements that form the point cloud by incorporating physical, chemical, and biological information relevant to the target problem.
+In this project, since solvation energy and Coulomb energy are closely related to the amino acid chains, we use two collections of point clouds.
+One consists of all carbon atoms, which form the backbone of amino acids.
+The other consists of all heavy atoms {C, N, O, S}, whose locations and interconnections are also critical to protein structures.
+Once the collections of point clouds are determined, the corresponding barcodes Ghrist:2007, which encode the persistent homology of the point clouds, are constructed.
+We use the GUDHI software package Boissonnat:2014 to generate these barcodes.
+The resulting barcodes are then processed by the algorithm below to produce vectors of topological features.
+Figure 2: The barcode of $H_{0},H_{1},H_{2}$ generated by the collection of cloud points using heavy atoms in proteins (a) 1a1e and (b) 1pb9.
+Figure 2 shows the barcodes of $H_{0},H_{1},H_{2}$ for protein 1a1e, the C-SRC Tyrosine Kinase, and protein 1pb9, the N-methyl-D aspartate receptor, using heavy atoms.
 The persistent homology analysis conducted in this study characterizes key biophysical phenomena, including intermolecular interactions, solvation effects, and hydrophobicity.
-Topological features are extracted from Betti-1 and Betti-2 barcodes, where the birth, death, and persistence values are computed based on the spatial distribution of carbon atoms and all heavy atoms within the protein structure, resulting in a total of 12 feature channels.
-Features derived from all heavy atoms represent the geometric interactions between specific types of atoms in the protein, whereas those computed from carbon atoms primarily reflect hydrophobic interactions and indirectly capture solvation effects.
-The analysis is performed over a distance scale of 0 to 50Å, which is discretized into uniform bins of 0.50Å to ensure fine-grained resolution of topological patterns.
-The topological information included in the barcode can be converted to features with the following approach [19].
+Topological features are extracted from $H_{1}$ and $H_{2}$ barcodes, where the birth, death, and persistence values are computed based on the spatial distribution of carbon atoms and all heavy atoms within the protein structure, resulting in a total of 12 feature channels.
+The significance of 1- and 2-dimensional homology groups lies in the fact that they represent 2-d rings and 3-d voids, respectively, which persist over a wide range of scales and are both topologically and geometrically meaningful.
+In contrast, the 0-dimensional homology group represents connected components, whose number decreases from the total number of points to 1 and therefore carries limited persistent signal.
+Higher-dimensional homology groups, although mathematically defined, are typically short-lived in persistence, behave as combinatorial noise, and lack clear geometric meaning.
+We use both the collection of heavy atoms and the collection of carbon atoms to capture critical structural and chemical information of proteins.
+Features derived from all heavy atoms represent geometric interactions between specific types of atoms in the protein, whereas those computed from carbon atoms primarily reflect hydrophobic interactions and indirectly capture solvation effects.
+The analysis is performed over a distance scale of 0 to 50Å, which is discretized into uniform bins of 0.25Å  to ensure fine-grained resolution of topological patterns.
+The topological information encoded in the barcodes is then converted into features using the following approach Cang:2017.
 We define the collection of barcodes as $\mathbb{B}(\alpha,\mathcal{C},\mathcal{D})$ with the following specifications.
@@
 $\mathcal{C}$ : type of simplicial complex (i.e., Ribs or Cech)
-$\mathcal{D}$ : dimension (i.e., Betti-0, Betti-1, etc.)
+$\mathcal{D}$ : dimension (i.e., $H_{1}$, $H_{2}$, etc.)
 Using the collection, the structured vectors ${\bf V}^{b}$, ${\bf V}^{d}$, and ${\bf V}^{p}$ can be constructed to respectively describe the birth, death, and persistent patterns of the barcodes in various spatial dimensions.
 Practically, the filtration interval $[0,L]$ is divided into $n$ equal length subintervals and the patterns are characterized on each subinterval.
-The description vectors using their $i$th component are defined as:
+The description vectors using their $i$th component for $1\leq i<n$, are defined as:
 […]
@@
 […]
+2.3.5 Feature generation from Persistent Image
+The topological features can also be constructed using persistence image alternatively.
+To this end, we first extract persistence diagrams from protein structures and convert them into stable, fixed-size image representations.
+The persistence diagrams are obtained separately from two types of point clouds: the collection of all carbon atoms and the collection of all heavy atoms.
+These two approaches are also used in the barcode-based feature construction as detailed previously.
+For each approach, we consider topological features in homology groups $H_{1}$ and $H_{2}$.
+Each persistence diagram is filtered using physically meaningful thresholds such that birth and persistence values are restricted to maximum bounds of $12.5$ Å.
+The resulting persistence domain is discretized using a resolution of $0.25$ Å, which produces a uniform image grid $D=[0,b_{\text{thr}}]\times[0,p_{\text{thr}}]$ with $b_{\text{thr}}=p_{\text{thr}}=50$ for all proteins.
+To emphasize the contribution of long-lived topological features, each point $(b_{i},p_{i})\in D$ is assigned a smooth exponential weight
+$\omega(b_{i},p_{i})=e^{p_{i}}-1.$ (11)
+(a) 1a0t (b) 1a4q
+Figure 3: Barcode (top) , persistence diagrams (middle) and persistence image (bottom) for proteins 1a0t (a) and 1a4q (b).
+For each protein, from left to right are: $H_{1}$ and $H_{2}$ for carbon atoms, $H_{1}$ and $H_{2}$ for all heavy atoms.
+This exponential weighting enhances features with larger persistence while assigning near-zero weight to short-lived topological noise, thereby improving the contrast between significant and insignificant structures in the resulting persistence images.
+In this project, persistence images are generated using the Gaussian kernel with bandwidth of $0.4$ Å, which is implemented by using the GUDHI library Boissonnat:2014.
+For each protein, four persistence images are produced: carbon atoms only for $H_{1}$ and $H_{2}$, all heavy atoms for $H_{1}$ and $H_{2}$.
+These four images are stacked to form a four-channel topological feature tensor, $\mathbf{X}\in\mathbb{R}^{b_{\text{thr}}\times p_{\text{thr}}\times 4}$.
+Finally, the resulting persistence image tensor is stored in the NumPy format and used as the input topological representation for downstream deep learning models.
+This approach provides a stable, compact, and fully vectorized encoding of protein topology thus is well suited for ML-based prediction.
+In Fig.
+3, the barcode of $H_{1}$ and $H_{2}$ using carbon atoms and using all heavy atoms for protein 1a0t and 1a4q are provided for comparison purpose with the persistent diagram and persist image.
+The persistence diagrams of the two proteins 1a0t and 1a4q show that most of the topological features are short-lived and concentrated near the diagonal, while only a small number of them persist.
+The corresponding persistence images highlight these persistent structures as bright and localized regions, with background noise effectively suppressed.
+The observed differences between persistence images from the cloud of points of carbon atoms and those from all heavy atoms show that topological invariants are influenced by the underlying chemical composition.
 2.4 Electrostatic Features
 2.4.1 Introduction
-In most of the molecular simulations, including Monte Carlo simulation, Brownian Dynamics, Molecular Dynamics, etc., the electrostatic interactions are characterized by the partial charges assigned at the atomic centers.
+Figure 4: An 2-d illustration for 3-d uniform and multi-scale electrostatic features for a protein (purple dashed line) with charges (black dots); Charges $q$ and reaction potential $\phi_{\text{reac}}$ are redistributed as point-multipoles (explosion symbols) using Cartesian treecode Li:2009 or FMM Tausch:2003 at the centers of the cluster at different levels (level 0: black; level 1: red; level 2: green).
+In most molecular simulations, including Monte Carlo simulation, Brownian Dynamics, Molecular Dynamics, etc., the electrostatic interactions are characterized by the partial charges assigned at the atomic centers.
 These partial charges are assigned according to the force fields, which are a set of mathematical equations and parameters used to calculate the potential energy of a molecule based on its atomic structure and interactions.
 The force fields are determined by experiment or quantum chemistry.
-Figure 3: An 2-d illustration for 3-d uniform and multi-scale electrostatic features for a protein (purple dashed line) with charges (shown as black dots); left: charges $q$ and reaction potential $\phi_{\text{reac}}$ are redistributed as point-multipoles (shown as explosion symbols) using Cartesian treecode [45] or FMM [46] at the centers of the cluster at different levels (level 0: black; level 1: red; level 2: green); right: charges $q$ and reaction potential $\phi_{\text{reac}}$ are redistributed at Chebyshev nodes (shown as square: level 0: black; level 1: red; level 2: green; only one cluster is shown in each level) using Barycentric Lagrangian treecode [56].
-The algorithm for obtaining the mulit-scale, physics-informed, uniform electrostatic features can be explained using the left of Fig.
-3.
+The algorithm for obtaining the mulit-scale, physics-informed, uniform electrostatic features can be explained using Fig.
+4.
 The electrostatic profile of the protein is determined by the partial charges $q_{n}({\bf r}_{n})$ or reaction potential $\phi_{\text{reac}}({\bf r}_{n})$ at the atomic centers ${\bf r}_{n}$’s as illustrated by the black dots for $n=1,2,\cdots,N_{c}$ with $N_{c}$ as the number of atoms of the protein.
@@
 The users specify $L$ and $p$ as needed.
-In the left of Fig.
-3.
+In Fig.
+4.
 the $N_{d}$ point multipoles (explosion symbols) are organized in a hierarchical way such that different level of details are given by different number of symbols at different levels.
@@
 Note when $|{\bf k}|=0$, ${\bf k}=(0,0,0)$; when $|{\bf k}|=1$, ${\bf k}\in\{(1,0,0),(0,1,0),(1,0,0)\}$; when $|{\bf k}|=2$, ${\bf k}\in\{(1,1,0),(1,0,1),(0,1,1),(2,0,0),(0,2,0),(0,0,2)\}$, etc.
-If the Cartesian treecode algorithm [45] is used, the computational cost is $O(N_{c}\log N_{c})$ for obtaining all the moments as detailed in [45], where $N_{c}$ is number of the particles/charges and $\log N_{c}$ is the level of trees.
-In fact, the computational cost can be as low as $O(N_{c})$ using the strategies in [46] such that the moments at the finest cluster are computed first and a M2M (moments to moments) transformation can be used to efficiently compute moments at any desired level from the bottom to the top.
+If the Cartesian treecode algorithm Li:2009 is used, the computational cost is $O(N_{c}\log N_{c})$ for obtaining all the moments as detailed in Li:2009, where $N_{c}$ is number of the particles/charges and $\log N_{c}$ is the level of trees.
+In fact, the computational cost can be as low as $O(N_{c})$ using the strategies in Tausch:2003 such that the moments at the finest cluster are computed first and a M2M (moments to moments) transformation can be used to efficiently compute moments at any desired level from the bottom to the top.
 Our simulations so far showed that the $O(N_{c}\log N_{c})$ cost is sufficiently fast for moments computation since $N_{c}$ is within the scale of tens of thousands even for large proteins.
@@
 The number of features up to level $L$ cluster (the largest box containing all particles is named as the level 0 cluster and each subdivision increases the level by 1) is
-$N_{f}(p,L)=N_{p}(1+8+8^{2}+\cdots+8^{L})=N_{p}\frac{8^{L+1}-1}{7}$
+$N_{f}(p,L)=N_{p}(1+8+8^{2}+\cdots+8^{L})=N_{p}\frac{8^{L+1}-1}{7}$ (13)
 where $N_{p}$ is the number of terms in multipole expansion with $N_{p}=1,4,10,20,35,56\cdots$ as the sequence of tetrahedral numbers when $p$th order multipole is used.
-Alternatively, the Barycentric treecode approach as shown in the right picture of Fig.
-3 can be used to generate the electrostatic features located the Chebyshev points (the squares) with modified weights as feature values.
-The details can be found in [56].
-For this approach, the computational cost is $O(N_{c}\log N_{c})$, and the number of terms in the similar sense of $N_{p}$ is $O(c_{p}^{3})$ where $c_{p}$ is the number of Chebyshev nodes in one direction.
-In this project, we focus on the Cartesian treecode approach, while the Barycentric treecode approach will be investigated in the future.
 2.4.2 Implementation Details
@@
 the PDB file) is obtained from publicly available repo e.g.
-the protein data bank [5].
+the protein data bank Berman:2000.
 The protein is represented by its atomic locations and partial charges: $q_{i}({\bf r}_{i})$ for $i=1,\cdots,N_{c}$, e.g.
-the PQR file using the PDB2PQR software [57].
+the PQR file using the PDB2PQR software Dolinsky:2007.
 Step 2.
@@
 The number of features up to level $L$ cluster is
-$N_{f}(p,L)=\frac{p(p+1)(p+2)(8^{L+1}-1)}{42}.$
+$N_{f}(p,L)=\frac{p(p+1)(p+2)(8^{L+1}-1)}{42}.$ (14)
 These $N_{f}$ numbers are ordered by level from 0 to $L$ and the coordinates of cluster centers in each level. This determines the dimension of our final feature vector $F\in\mathbb{R}^{N_{f}}$.
@@
 Step 3.
-The Cartesian treecode [45] will be called and the moments for the multipoles will be output as the electrostatic features.
-These features can be used as part of the inputs for machine learning algorithms, which will be shown the next.
-2.5 A DNN based machine learning model with topological and electrostatic features
-Figure 4: The DNN based Machine Learning model which uses protein structural data, force field, and known protein properties to mathematically generate electrostatic, topological, algebraic, geometric, and features to train a learned model and then use the learned model to predict unknown properties for protein with available structures.
+The Cartesian treecode Li:2009 will be called and the moments for the multipoles will be output as the electrostatic features.
+These features can be used as part of the inputs for ML algorithms, which will be shown next.
+2.5 The Machine Learning Model
+Figure 5: The DNN based Machine Learning model using protein structures, force field, and known protein properties to mathematically generate electrostatic, topological, algebraic, geometric, and features to train a learned model and then use the learned model to predict unknown properties for protein with available structures.
+This project focuses on electrostatic and topological features.
 The main goal of this project is to discover the hidden and useful information embedded in the protein structural data from the protein data bank in a simple and abstract way.
@@
 In this project, we put our focus on the electrostatic and topological features.
-The DNN based Machine Learning model is shown in Fig.
-4, which uses the available protein structural data [58], force field such as AMBER [59], CHARMM[60], AMOEBA [61], etc.
-and known protein properties repositories such as PDBbind Database [62], Protein pKa Database [63], etc to mathematically generate features to train a learned model and then use the learned model to predict unknown properties for protein with available structures.
-In this project, we focus on the effectiveness of topological and electrostatic features using computed Coulombic energies and solvation free energies as labels.
-3 Results
-3.1 Specification of the machine learning model
-We provide more details of the machine learning model with focus on the neural network and its related specification.
-The predictive model developed in this study is designed to predict solvation and Coulomb energies based on a combination of topological and electrostatic features.
-Prior to model training, outliers are identified and removed using the interquartile range (IQR) method.
-All input features and target labels are standardized to have zero mean and unit variance using the StandardScaler() utility from sklearn.preprocessing.
-The inverse transformations are applied to the output, returning values with physical meanings e.g.
-energies in kcal/mol during post-processing for evaluation purposes.
-The network architecture of the model as shown in Fig.
-5 is implemented using the Keras Sequential API with two distinct branches.
-The first branch as color-coded in blue is a one-dimensional convolutional neural network (CNN) processing the topological features.
-The CNN branch has three convolution-pooling layers with (128, 64, 64) filters and ($\tanh$, relu, $\tanh$) activation function respectively.
-The 3 by 3 kernels are used.
-This CNN branch is responsible for capturing hierarchical and spatially correlated patterns in the topological data.
-The output of the CNN is flattened and passed through a subsequent Dense layer of 64 units.
-The other branch is the electrostatic features into two fully connected hidden dense layers with (128, 64) units respectively as shown in orange.
-After independent processing, the outputs from both branches are concatenated and passed through additional three hidden dense layers with (32, 16, 8) units as shown in purple to produce the final regression output of energy as shown in red.
+The DNN based ML model is shown in Fig.
+5, which uses the available protein structural data PDB, force field such as AMBER Perlman:1995, CHARMMBrooks:1983, AMOEBA Zhang:2018, etc.
+and known protein properties repositories such as PDBbind Database Su:2019, Protein pKa Database Pahari:2019, etc to mathematically generate features to train a learned model and then use the learned model to predict unknown properties for protein with available structures.
+In this project, we focus on the effectiveness of topological and electrostatic features using computed Coulomb energies and solvation energies as labels.
+Mathematically, the surrogate models (Coulomb energies computed by treecode or solvation energies from the numerically solved PB model)
+$E=f(\mathbf{X},\mathbf{Q})$ (15)
+are approximated by the machine learning models
+$\hat{E}=\hat{f}(F_{e},F_{t},\Theta)$ (16)
+where $E$ and $\hat{E}$ are the energies, $\mathbf{X}$ and $\mathbf{Q}$ are the positions and charges of protein atoms, $\Theta$ are parameters in the ML model to be trained by data, $F_{e}=F_{e}(\mathbf{X},\mathbf{Q})$ and $F_{t}=F_{t}(\mathbf{X})$ are electrostatic features and topological features and their generators, which are non-trivial mappings from protein structure and electrostatics to features.
+2.6 Specification of the ML Model
+The model developed in this study is to to predict Coulomb and solvation energies using combined topological and electrostatic features.
+All target labels were standardized to zero mean and unit variance, and predictions were inverse-transformed to recover physical units (kcal/mol) for evaluation.
+As illustrated in Fig.
+6, the proposed architecture consists of two branches.
+The first branch, shown in blue, is a one-dimensional convolutional neural network (CNN) designed to process the topological features.
+These features are passed through three consecutive convolution–pooling blocks (with 128, 64, and 64 filters; kernel size = 3; and activation functions tanh, ReLU, and tanh, respectively), followed by a fully connected layer with 64 units.
+The electrostatic features are processed through two fully connected hidden layers with 128 and 64 units, respectively (shown in orange).
+After independent processing in these two branches, their outputs are concatenated and passed through three additional fully connected layers with 32, 16, and 8 units (shown in purple).
+The final dense layer then produces the regression output for the predicted energy (shown in red).
+To improve generalization, dropout, $l_{2}$ regularization, and batch normalization were applied.
+The model was trained using the Adam optimizer (learning rate $1.0\times 10^{-4}$) with mean squared error as the loss function.
+The dataset was randomly divided into 80% training and 20% testing subsets.
+Model selection was performed using five-fold cross-validation on the training set, and the best-validated model was subsequently evaluated on the held-out test set.
+The model architecture and hyperparameters were selected based on empirical considerations and adopted from previously published studies Cang:2017.
+The complete hyperparameter configuration is provided in Table $S1$ of the Supporting Information.
 Figure 6: The architecture of the deep neural network with two branches taking in topological features and electrostatic features.
 The outputs of the two branches are concatenated into one deep neural network.
-To enhance generalization and mitigate overfitting, batch normalization is applied after intermediate layers, and dropout is employed to randomly deactivate neurons during training.
-The model is trained for 500 epochs using the Adam optimizer with a learning rate of 0.0001 and a batch size of 32.
-The mean squared error (MSE) is used as a loss function.
-Model architecture and hyperparameters are selected through empirical tuning and informed by previous studies employing Hyperopt [19, 64].
-The dataset as described below was split randomly into 80% for training and 20% for testing, and the model was trained with 5-fold cross-validation as well.
 2.7 Data
-The datasets are sourced from the PDBbind database [65].
-In particular, we categorize used data into two datasets as described below.
-The Amber ff14SB general force field is applied for the atomic van der Waals radii and partial charges when the PDB2PQR [57] software is used.
-Dataset 1: These data are from the refined set of the v2018 release [65].
-An interquartile range method is used to remove outliers based on $E_{\text{coul}}$, resulting in 4,658 complexes used in our simulation.
-We also we computed the electrostatic solvation free energies $E_{\text{solv}}$ of the proteins, as part of the 4,658 protein-ligand complexes.
-All the $E_{\text{solv}}$ values are computed using the 2nd order accurate Matched Interface and Boundary Poisson-Boltzmann (MIBPB) solver [20] at mesh size h=0.2h=0.2Å.
-Dataset 2: These data are from the the v2020 release, which has a general set with 14,127 complexes and refined set with 5,316 complexes.
-We computed the Coulomb energies $E_{\text{coul}}$ of the all proteins using the Treecode [45].
-An interquartile range method is used to remove outliers based on $E_{\text{coul}}$, resulting in 17,007 complexes used in our simulation.
+The datasets used in this study are sourced from the PDBbind database Liu:2014, a widely used benchmark for protein–ligand complexes.
+We categorize the data into two datasets as described below.
+For all structures, atomic van der Waals radii and partial charges are assigned using the AMBER ff14SB force field through the PDB2PQR software Dolinsky:2007.
+[…]
+Dataset 2.
+This dataset is constructed from the PDBbind v2020 release, comprising a general set of 14,127 complexes and a refined set of 5,316 complexes.
+Coulombic energies $E_{\mathrm{coul}}$ are computed for all proteins using a Cartesian treecode algorithm Li:2009.
+To ensure data quality, outliers are identified and removed using the IQR method, applying thresholds of [1,164.23, 338.04][-1,164.23,\;338.04] kcal/mol, where $Q_{1}=-600.88$ kcal/mol,  $Q_{3}=-225.31$ kcal/mol, $\text{IQR}=375.57$ kcal/mol.
+Following this filtering step, a total of 17,007 protein structures are retained for subsequent model training and evaluation.
+(a)                (b)                (c)
+Figure 7: Box-and-whisker plots showing the distributions of Coulombic energy $E_{\mathrm{coul}}$ for Dataset 1 in (a) and Dataset 2 in (b), and solvation energy $E_{\mathrm{solv}}$ for Dataset 1 in (c).
+Jittered points represent individual protein samples.
+Figure 7 illustrates the distributions of Coulombic and solvation energies across the two datasets.
+The Coulombic energies in both datasets are concentrated within a relatively narrow negative range, whereas the solvation energies exhibit a substantially broader distribution with greater dispersion.
+This contrast reflects the fundamentally different physical scales and variability associated with Coulombic interactions and solvation effects.
 2.8 Research Purposes and Metrics
-We here provide some simulation results to demonstrate the performance of our model. These results serve three research purposes.
-Purpose 1: Training on larger data set receives better performance.
-Purpose 2: When electrostatic features are considered, performance of model improves when when treecode level $L$, Taylor expansion order $p$, and their associated number of features $N_{f}$ increases.
-Purpose 3: Using both electrostatic features and topological features outperforms using only one kind of these features.
-Metrics: We use three common metrics to measure the performance of the model as defined below.
+We provide simulation results to evaluate the performance of the proposed model. These results are designed to address the following three research purposes:
+Purpose 1: Training on a larger dataset leads to improved predictive performance.
+Purpose 2: When electrostatic features are incorporated, model performance improves as the treecode level $L$, Taylor expansion order $p$, and the corresponding number of features $N_{f}$ increase.
+Purpose 3: Combining electrostatic and topological features outperforms models that use only one type of feature.
+Metrics: We use three common metrics to measure the performance of the model, as defined below.
 In all these metrics, we use $y_{i}$ to denote actual value, $\hat{y}_{i}$ for predicted value, $\bar{y}$ for mean of actual values, $\bar{\hat{y}}$ for mean of predicted values.
-MAPE: The mean average percentage errors, which measures prediction accuracy as a percentage. Not suitable if the actual values contain zero or values close to zero.
-$\text{MAPE}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|\times 100\%$
-MSE: The mean-squared error, which measures the average absolute difference between predicted and actual values.
-$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$
-$R^{2}$, which indicates the proportion of variance in the dependent variable that is predictable from the independent variables.
-$R^{2}=1-\frac{\text{SS}_{\text{res}}}{\text{SS}_{\text{tot}}}$
+MAPE: The mean average percentage errors, which measures prediction accuracy as a percentage. This metric is not suitable if the actual values contain zero or values close to zero.
+$\text{MAPE}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|\times 100\%$ (17)
+MSE: The mean-squared error, which measures the average squared difference between predicted and actual values.
+$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (18)
+$R^{2}$, the quantity indicating the proportion of variance in the dependent variable that is predictable from the independent variables.
+$R^{2}=1-\frac{\text{SS}_{\text{res}}}{\text{SS}_{\text{tot}}}$ (19)
 where $\text{SS}_{\text{tot}}=\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$ measures the total variance in the observed data and $\text{SS}_{\text{res}}=\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ measures the variance not explained by the model.
-3.4 Performance
-We next provide simulation results using different dataset and models.
-On reporting and explaining the performance of each combination of model and dataset, we focus on the achievement of the three aforementioned purposes.
-$E_{\text{coul}}$ computed using electrostatic features only on Dataset 1 and Dataset 2
-We first investigate the effectiveness of electrostatic features in the prediction of the Coulombic energy $E_{\text{coul}}$.
-Note $E_{\text{coul}}$ is determined by the location and charges of the particles, which also determines the moments of the multipoles as the electrostatic features.
-Table 1: $E_{\text{coul}}$ predicted using Electrostatic Features only on Dataset 1 and Dataset 2; D2-D1 indicates the difference of corresponding values of MSE and $R^{2}$ from Dataset 1 and Dataset 2.
-$p$ $L$ $N_{f}(p,L)$ Dataset 1 Dataset 2 D2-D1
+3 Results
+We next provide simulation results using different datasets and models.
+Following that, additional comparisons with baseline models and different grouping strategies are provided to justify the learned model.
+3.1 Model Performance
+In reporting and explaining the performance of each combination of model and dataset, we focus on achieving the three aforementioned purposes.
+[…]
+3.1.1 $E_{\text{coul}}$ predicted using electrostatic features only on Dataset 1 and Dataset 2
+We first investigate the effectiveness of electrostatic features in predicting the Coulombic energy $E_{\text{coul}}$.
+Note that $E_{\text{coul}}$ is determined by the locations and charges of the particles, which also determine the moments of the multipoles as electrostatic features.
+Table 1: $E_{\text{coul}}$ (kcal/mol) prediction results using electrostatic features only on Dataset 1 and Dataset 2.
+The D2–D1 columns report the differences (Dataset 2 - Dataset 1) for the corresponding metrics.
+$p$ $L$ $N_{f}(p,L)$ Dataset 1 Dataset 2 D2–D1
 MSE MAPE $R^{2}$ MSE MAPE $R^{2}$ MSE MAPE $R^{2}$
-1 0 4 0.641 0.328 0.351 0.682 0.532 0.331 0.041 0.204 -0.020
-2 0 10 0.426 0.253 0.569 0.351 0.352 0.655 -0.074 0.100 0.086
-3 0 20 0.352 0.206 0.643 0.252 0.318 0.753 -0.100 0.111 0.110
-4 0 35 0.279 0.190 0.717 0.270 0.361 0.735 -0.009 0.172 0.018
-0 1 9 0.699 0.354 0.292 0.663 0.548 0.349 -0.036 0.194 0.057
-1 1 36 0.295 0.206 0.701 0.235 0.277 0.769 -0.060 0.070 0.068
-2 1 90 0.284 0.175 0.712 0.154 0.224 0.848 -0.129 0.049 0.136
-3 1 180 0.207 0.178 0.791 0.197 0.260 0.806 -0.009 0.082 0.016
-4 1 315 0.203 0.157 0.794 0.224 0.320 0.780 0.021 0.163 -0.014
-0 2 73 0.469 0.261 0.525 0.337 0.332 0.669 -0.131 0.071 0.144
-1 2 292 0.249 0.182 0.748 0.144 0.201 0.859 -0.105 0.019 0.111
-2 2 730 0.227 0.177 0.770 0.170 0.217 0.833 -0.057 0.040 0.063
-3 2 1460 0.211 0.166 0.786 0.162 0.220 0.841 -0.049 0.054 0.055
-4 2 2555 0.219 0.183 0.778 0.180 0.297 0.823 -0.039 0.114 0.045
-0 3 585 0.465 0.265 0.529 0.287 0.283 0.719 -0.179 0.018 0.190
-1 3 2340 0.513 0.220 0.480 0.255 0.259 0.750 -0.259 0.039 0.270
-2 3 5850 0.469 0.230 0.525 0.215 0.240 0.789 -0.254 0.010 0.264
-3 3 11700 0.450 0.220 0.544 0.269 0.248 0.736 -0.181 0.028 0.192
-4 3 20475 0.451 0.221 0.543 0.241 0.323 0.764 -0.210 0.102 0.221
+1 0 4 0.584 0.337 0.385 0.682 0.531 0.323 0.098 0.194 -0.062
+2 0 10 0.353 0.246 0.628 0.340 0.331 0.663 -0.013 0.085 0.035
+3 0 20 0.254 0.207 0.733 0.236 0.284 0.767 -0.017 0.077 0.035
+4 0 35 0.325 0.186 0.668 0.249 0.321 0.753 -0.076 0.135 0.085
+0 1 9 0.708 0.344 0.254 0.633 0.498 0.372 -0.075 0.154 0.118
+1 1 36 0.256 0.205 0.730 0.241 0.273 0.761 -0.015 0.068 0.030
+2 1 90 0.219 0.177 0.769 0.156 0.230 0.845 -0.063 0.053 0.076
+3 1 180 0.169 0.155 0.822 0.163 0.237 0.838 -0.006 0.083 0.016
+4 1 315 0.185 0.173 0.805 0.165 0.234 0.837 -0.020 0.061 0.032
+0 2 73 0.403 0.278 0.575 0.322 0.324 0.681 -0.081 0.047 0.105
+1 2 292 0.234 0.223 0.754 0.127 0.206 0.874 -0.106 -0.017 0.120
+2 2 730 0.204 0.187 0.786 0.133 0.208 0.868 -0.070 0.021 0.082
+3 2 1460 0.182 0.163 0.808 0.123 0.195 0.880 -0.059 0.032 0.071
+4 2 2555 0.247 0.225 0.740 0.145 0.229 0.857 -0.102 0.004 0.118
+0 3 585 0.393 0.239 0.586 0.265 0.279 0.737 -0.128 0.041 0.151
+1 3 2340 0.443 0.211 0.533 0.253 0.236 0.752 -0.190 0.024 0.219
+2 3 5850 0.411 0.206 0.567 0.217 0.252 0.784 -0.194 0.046 0.218
+3 3 11700 0.472 0.249 0.503 0.209 0.261 0.792 -0.262 0.012 0.290
+4 3 20475 0.381 0.240 0.598 0.223 0.267 0.779 -0.158 0.027 0.180
 (a)                 (b)                 (c)
-Figure 6: (a) $R^{2}$ of model model performance against number of features; (b) scatter plot for Dataset 1 ($p=4$, $L=1$ with total 315 features); (c) scatter plot for Dataset 2 ($p=1$, $L=2$ with total 292 features)
-Table 1 reports the $E_{\text{coul}}$ prediction using electrostatic only on Dataset 1 and Dataset 2.
-The first important pattern we can observe from the table is that the MSE and $R^{2}$ obtained from Dataset 2 are essentially better than those obtained from Dataset 1 as the differences of MSE (D2-D1) are almost all negative and the difference of $R^{2}$ (D2-D1) is almost all positive.
-This pattern validates Purpose 2 indicating advantage of using larger dataset, particularly at capturing more overall variance in the target.
-We also noticed the difference of MAPE(D2-D1) is almost all positive, which apparently contradicts the results from MSE and $R^{2}$.
-In fact, this gives more information about the data distribution, i.e.
-the data in Dataset 2 but not in Dataset 1 might contain more smaller target values, yielding large percentage errors.
-Also from Table 1, we can see on Dataset 1, MSE decreases and $R^{2}$ increase with the increment of $p$ and $L$, reaching optimized value at $p=4$ and $L=1$ with 315 features.
-Similarly on Dataset 2, MSE decreases and $R^{2}$ increases with the increment of $p$ and $L$ too, reaching optimized value at $p=1$ and $L=2$ with 292 features.
-This pattern validates the Purpose 1 showing the advantage of using more features.
-However, we should aware that optimized $p$ and $L$ are affected by the size of the training dataset.
-When the number of features are larger than number of samples, increasing number of features might bring negative impact such as overfitting.
-We plot $R^{2}$ in Fig.
-6(a) with horizontal axis as number of features $N_{f}$ from the third column of Table 1.
-The validation of Purpose 1 is shown since blue squares are lower than the red ones.
-The validation of Purpose 2 is also shown since markers of the same color contain an increasing pattern as number of features increase until there are too many features.
-We also plot the best cases on using both Datasets in Fig.
-6(b,c) and we can see the distribution of the data points are essentially around the $\hat{y}=y$ line.
-$E_{\text{coul}}$ computed using e-features and both features on Dataset 2
-Table 2: $E_{\text{coul}}$ computed using e-features and both features on Dataset 2; the MSE and $R^{2}$ for using Topological Features only is 0.039 and 0.962 respectively.
-$p$ $L$ $N(p,L)$ e-features both features difference
-MSE $R^{2}$ MSE $R^{2}$ MSE $R^{2}$
-1 0 4 0.682 0.331 0.039 0.961 -0.643 0.631
-2 0 10 0.351 0.655 0.033 0.967 -0.318 0.312
-3 0 20 0.252 0.753 0.044 0.956 -0.208 0.203
-4 0 35 0.270 0.735 0.030 0.970 -0.240 0.235
-0 1 9 0.663 0.349 0.030 0.970 -0.634 0.621
-1 1 36 0.235 0.769 0.040 0.960 -0.195 0.191
-2 1 90 0.154 0.848 0.044 0.956 -0.110 0.108
-3 1 180 0.197 0.806 0.037 0.963 -0.160 0.156
-4 1 315 0.224 0.780 0.045 0.955 -0.179 0.175
-0 2 73 0.337 0.669 0.048 0.952 -0.289 0.283
-1 2 292 0.144 0.859 0.062 0.939 -0.082 0.080
-2 2 730 0.170 0.833 0.051 0.949 -0.119 0.116
-3 2 1460 0.162 0.841 0.032 0.969 -0.131 0.128
-4 2 2555 0.180 0.823 0.039 0.961 -0.141 0.138
-0 3 585 0.287 0.719 0.047 0.953 -0.240 0.235
-1 3 2340 0.255 0.750 0.050 0.951 -0.205 0.201
-2 3 5850 0.215 0.789 0.042 0.958 -0.172 0.169
-3 3 11700 0.269 0.736 0.032 0.968 -0.236 0.232
-4 3 20475 0.241 0.764 0.037 0.964 -0.204 0.200
-We next investigate the effectiveness of combined features.
-This investigation has two parts.
-The first part is using $E_{\text{coul}}$ as labels with Dataset 2 and the second part is using $E_{\text{solv}}$ as labels with Dateset 1 considering the availability of labels and the size of the datasets.
-The MSE and $R^{2}$ for different models are shown in Table 2.
-Note the MSE and $R^{2}$ for using topological features only is 0.039 and 0.962 respectively, which is already very effective.
-If we see the difference column, we can see that adding topological features fundamentally improves the model using electrostatic features only by decreasing MSE and increase $R^{2}$.
-This partially validates purpose 2.
-Note we did not report MAPE result here as it shows similar pattern as MSE does.
-If we consider topological features first, adding electrostatic features seems only occasionally improve the results.
-This is because the results from using topological features are too good for the effectiveness of electrostatic features to kicked in.
-Our next test case shows when the topological features are relatively less effective, adding the electrostatic features can significantly improve the performance.
-Before we move to that case, let’s linger on Fig.
-7, which is the scatter plot of predicted values vs true values for using topological features, electrostatic feature and combined feature, for the observation that the improvement from (b) to (c) are very obvious while the improvement from (a) to (c), although not that significant, is still discernible.
+Figure 8: (a) $R^{2}$ of model performance for models trained on the two datasets against number of features; (b) scatter plot for Dataset 1 ($p=3$, $L=1$ with total 180 features); (c) scatter plot for Dataset 2 ($p=3$, $L=2$ with total 1460 features)
+Table 1 reports the $E_{\text{coul}}$ predictions obtained using electrostatic features only on Dataset 1 and Dataset 2.
+The first important pattern observed from the table is that the MSE and $R^{2}$ obtained from Dataset 2 are generally better than those from Dataset 1, as the differences in MSE (D2-D1) are almost all negative, while the difference in $R^{2}$ (D2-D1) is almost all positive.
+This pattern validates Purpose 2, indicating the advantage of using a larger dataset, particularly in capturing more overall variance in the target.
+We also notice that the differences in MAPE(D2-D1) are almost all positive, apparently contradicting the trends observed for MSE and $R^{2}$.
+In fact, this behavior provides additional insight into the data distribution: Dataset 2, but not Dataset 1, contains more samples with smaller target values as shown in Fig.
+7, leading to larger percentage errors.
+From Table 1, we also observe that in Dataset 1, the MSE decreases and $R^{2}$ increases as $p$ and $L$ increase, reaching an optimal value at $p=3$ and $L=1$ with 180 features.
+Similarly, in Dataset 2, the MSE decreases and $R^{2}$ increases with increasing $p$ and $L$, reaching an optimal value at $p=3$ and $L=2$ with 1460 features.
+This pattern validates Purpose 1, demonstrating the advantage of using more features.
+However, we should be aware that the optimal choices of $p$ and $L$ are influenced by the size of the training dataset.
+When the number of features exceeds the number of samples, increasing the number of features may have a negative impact, such as overfitting.
+We plot the $R^{2}$ values in Fig.
+8(a), where the horizontal axis represents the number of features $N_{f}$ given in the third column of Table 1.
+The validation of Purpose 1 is evident, as the blue squares are consistently lower than the red ones.
+The validation of Purpose 2 is also observed, as markers of the same color exhibit an increasing trend with the number of features until the feature set becomes excessively large.
+We also plot the best-performing cases for both datasets in Fig.
+8(b) and (c), where the reference-predicted values are distributed closely around the $\hat{y}=y$ line.
+3.1.2 $E_{\text{coul}}$ predicted using e-features and both features on Dataset 2
+Table 2: $E_{\text{coul}}$ (kcal/mol) prediction results using electrostatic features and combined features on Dataset 2.
+The corresponding MSE, MAPE, and $R^{2}$ values obtained using Topological Features only are 0.030, 0.071, and 0.970, respectively.
+$p$ $L$ $N_{f}(p,L)$ e-features only both features difference
+MSE MAPE $R^{2}$ MSE MAPE $R^{2}$ MSE MAPE $R^{2}$
+1 0 4 0.682 0.531 0.323 0.028 0.076 0.972 -0.654 -0.454 0.649
+2 0 10 0.340 0.331 0.663 0.035 0.072 0.966 -0.305 -0.258 0.303
+3 0 20 0.236 0.284 0.767 0.031 0.064 0.969 -0.205 -0.220 0.202
+4 0 35 0.249 0.321 0.753 0.027 0.082 0.973 -0.222 -0.239 0.221
+0 1 9 0.633 0.498 0.372 0.029 0.067 0.971 -0.604 -0.431 0.599
+1 1 36 0.241 0.273 0.761 0.030 0.068 0.971 -0.211 -0.205 0.210
+2 1 90 0.156 0.230 0.845 0.029 0.063 0.971 -0.127 -0.167 0.126
+3 1 180 0.163 0.237 0.838 0.029 0.067 0.971 -0.134 -0.170 0.133
+4 1 315 0.165 0.234 0.837 0.024 0.073 0.976 -0.141 -0.162 0.140
+0 2 73 0.322 0.324 0.681 0.031 0.069 0.969 -0.291 -0.255 0.289
+1 2 292 0.127 0.206 0.874 0.035 0.131 0.965 -0.092 -0.075 0.092
+2 2 730 0.133 0.208 0.868 0.033 0.109 0.968 -0.101 -0.099 0.100
+3 2 1460 0.123 0.195 0.880 0.033 0.092 0.967 -0.089 -0.103 0.087
+4 2 2555 0.145 0.229 0.857 0.026 0.067 0.974 -0.119 -0.162 0.117
+0 3 585 0.265 0.279 0.737 0.031 0.071 0.969 -0.234 -0.209 0.232
+1 3 2340 0.253 0.236 0.752 0.034 0.077 0.967 -0.219 -0.159 0.215
+2 3 5850 0.217 0.252 0.784 0.031 0.075 0.969 -0.186 -0.177 0.185
+3 3 11700 0.209 0.261 0.792 0.029 0.083 0.971 -0.181 -0.178 0.179
+4 3 20475 0.223 0.267 0.779 0.023 0.074 0.978 -0.200 -0.193 0.199
+We next investigate the effectiveness of the combined feature representation.
+This analysis consists of two parts.
+In the first part, $E_{\text{coul}}$ is used as the label on Dataset 2, while in the second part, $E_{\text{solv}}$ is used as the label on Dataset 1, taking into account the size of the datasets and the availability of labels.
 (a)                 (b)                 (c)
-Figure 7: scatter plots: (a) topological features only, (b) e-features only ($p=1$, $L=2$ with total 292 features), (c) both features ($p=4$, $L=0$ with total 35 features).
-$E_{\text{solv}}$ computed using e-feature and both features on Dataset 1
+Figure 9: Scatter plots for the prediction of $E_{\text{coul}}$ on Dataset 2: (a) topological features only, (b) e-features only ($p=3$, $L=2$ with total 1460 features), (c) both features ($p=4$, $L=1$ with total 315 features).
+The MSE, MAPE, and $R^{2}$ values for the different models are reported in Table 2.
+Using topological features alone yields MSE, MAPE, and $R^{2}$ values of 0.030, 0.071, and 0.970, respectively, indicating strong predictive accuracy.
+Moreover, the difference column clearly shows that incorporating topological features into the electrostatics-only model consistently reduces MSE and MAPE while increasing $R^{2}$.
+These results partially support Purpose 2.
+If we consider topological features first, it appears that adding electrostatic features does not significantly improve the results.
+This is because the performance obtained using topological features is already very strong, leaving limited room for the effectiveness of electrostatic features to manifest.
+Our next test case shows that when the topological features are relatively less effective, adding electrostatic features can significantly improve performance.
+Before moving to that case, we briefly examine Fig.
+9, which shows scatter plots of predicted values versus true values using topological features, electrostatic features, and combined features.
+The improvement from (b) to (c) is very evident, while no obvious improvement is observed from (a) to (c).
+3.1.3 $E_{\text{solv}}$ predicted using e-features only and both features on Dataset 1
 As mentioned earlier, computing electrostatic solvation energies accurately is time consuming.
-Here we computed the solvation energies for the proteins from Dataset 1 using the MIBPB [20] solver.
-These solvation energies are used as labels for the neural network training and testing.
-A similar simulation using Dataset 1 are conducted and the results are reported in Table 3.
-Form the table, we can see that the improvement on using both features from using topological features only or from using electrostatic features are significant.
-If we focus on the difference column (both features - e-features only), we can see adding topological features completely improve the model using electrostatic features only by decreasing MSE and increase $R^{2}$.
-Meanwhile, the MSE and $R^{2}$ in the columns of using both features are obviously better than using topological feature only, which has MSE = 0.150 and $R^{2}$= 0.826.
-These results validate purpose 3, showing the advantage of using topological and electrostatic features combined.
-Again, we did not report MAPE result here as it shows similar pattern as MSE does.
-Table 3: $E_{\text{solv}}$ computed using e-features and both features on Dataset 1; the MSE and $R^{2}$ for using Topological Features only is 0.150 and 0.826 respectively.
-$p$ $L$ $N(p,L)$ e-features both features difference
-MSE $R^{2}$ MSE $R^{2}$ MSE $R^{2}$
-1 0 4 0.345 0.604 0.063 0.926 -0.282 0.323
-2 0 10 0.271 0.690 0.090 0.895 -0.180 0.206
-3 0 20 0.135 0.845 0.064 0.926 -0.071 0.081
-4 0 35 0.121 0.861 0.065 0.925 -0.056 0.064
-0 1 9 0.447 0.487 0.114 0.867 -0.333 0.380
-1 1 36 0.148 0.830 0.077 0.911 -0.072 0.081
-2 1 90 0.095 0.891 0.084 0.903 -0.012 0.012
-3 1 180 0.111 0.873 0.105 0.878 -0.006 0.005
-4 1 315 0.128 0.853 0.097 0.887 -0.031 0.034
-0 2 73 0.214 0.754 0.092 0.892 -0.122 0.138
-1 2 292 0.169 0.806 0.127 0.852 -0.041 0.045
-2 2 730 0.164 0.812 0.128 0.851 -0.036 0.039
-3 2 1460 0.212 0.757 0.138 0.839 -0.074 0.082
-4 2 2555 0.218 0.750 0.165 0.808 -0.053 0.058
-0 3 585 0.336 0.615 0.148 0.827 -0.188 0.213
-1 3 2340 0.399 0.542 0.218 0.747 -0.181 0.205
-2 3 5850 0.390 0.553 0.171 0.801 -0.219 0.249
-3 3 11700 0.469 0.461 0.195 0.773 -0.275 0.312
-4 3 20475 0.365 0.581 0.242 0.718 -0.123 0.137
-The optimized case for using e-feature only happens at $p=1$ and $L=1$ with 90 features.
-Adding topological features improve the performance.
-We provide the scatter plot between reference values and predicted values for using topological features only, electrostatic features only, and both features in Fig.
-8.
-The improvements of using two kind of features as shown in (c) can be seen as compared with topological features only in (a) and electrostatic features only in (b).
+Here, we use the MIBPB solver Chen:2011 to compute the solvation energies of proteins in Dataset 1, which are then used as labels for model training and testing.
+A similar simulation using Dataset 1 is conducted, and the results are reported in Table 3.
+From the table, we can see that the improvement obtained by using both features, compared to using only topological features or only electrostatic features, is significant.
+If we focus on the difference column, we see that adding topological features clearly improves the electrostatics-only model by reducing the MSE and MAPE while increasing the $R^{2}$ score.
+In addition, the results obtained using both features are noticeably better than those obtained using topological features alone, which yield MSE = 0.150, MAPE = 0.153, and $R^{2}$ = 0.826.
+These results validate purpose 3, and demonstrate the advantage of combining topological and electrostatic features.
+Table 3: $E_{\text{solv}}$(kcal/mol) computed using e-features and both features on Dataset 1; the MSE, MAPE and $R^{2}$ for using Topological Features only is 0.150, 0.153 and 0.826 respectively.
+$p$ $L$ $N_{f}(p,L)$ e-features only both features difference
+MSE MAPE $R^{2}$ MSE MAPE $R^{2}$ MSE MAPE $R^{2}$
+1 0 4 0.345 0.243 0.604 0.063 0.108 0.926 -0.282 -0.136 0.323
+2 0 10 0.271 0.256 0.690 0.090 0.094 0.895 -0.180 -0.162 0.206
+3 0 20 0.135 0.142 0.845 0.064 0.081 0.926 -0.071 -0.061 0.081
+4 0 35 0.121 0.138 0.861 0.065 0.083 0.925 -0.056 -0.055 0.064
+0 1 9 0.447 0.318 0.487 0.114 0.169 0.867 -0.333 -0.149 0.380
+1 1 36 0.148 0.161 0.830 0.077 0.104 0.911 -0.072 -0.057 0.081
+2 1 90 0.095 0.130 0.891 0.084 0.098 0.903 -0.012 -0.032 0.012
+3 1 180 0.111 0.138 0.873 0.105 0.103 0.878 -0.006 -0.036 0.005
+4 1 315 0.128 0.144 0.853 0.097 0.105 0.887 -0.031 -0.039 0.034
+0 2 73 0.214 0.205 0.754 0.092 0.116 0.892 -0.122 -0.089 0.138
+1 2 292 0.169 0.156 0.806 0.127 0.123 0.852 -0.041 -0.033 0.045
+2 2 730 0.164 0.156 0.812 0.128 0.116 0.851 -0.036 -0.039 0.039
+3 2 1460 0.212 0.168 0.757 0.138 0.124 0.839 -0.074 -0.045 0.082
+4 2 2555 0.218 0.168 0.750 0.165 0.128 0.808 -0.053 -0.040 0.058
+0 3 585 0.336 0.211 0.615 0.148 0.135 0.827 -0.188 -0.076 0.213
+1 3 2340 0.399 0.216 0.542 0.218 0.143 0.747 -0.181 -0.073 0.205
+2 3 5850 0.390 0.202 0.553 0.171 0.139 0.801 -0.219 -0.063 0.249
+3 3 11700 0.469 0.242 0.461 0.195 0.145 0.773 -0.275 -0.097 0.312
+4 3 20475 0.365 0.210 0.581 0.242 0.156 0.718 -0.123 -0.053 0.137
+The optimized case for the electrostatic-only model occurs at $p=2$ and $L=1$ with 90 features.
+Adding topological features further improves the performance, achieving an MSE of 0.064, a MAPE of 0.081, and an $R^{2}$ value of 0.926 using only 20 features at $p=3$ and $L=0$.
+Figure 10 presents scatter plots for the optimized cases, comparing reference and predicted values for models using topological features only, electrostatic features only, and the combined feature set.
+The benefits of combining both types of features, as illustrated in (c), can be clearly observed when compared with using only topological features in (a) and only electrostatic features in (b).
 (a)                 (b)                 (c)
-Figure 8: scatter plots: (a) topological features only, (b) e-features only ($p=2$, $L=1$ with total 90 features), (c) both features.
-4 Software Dissemination
-The code associated with this project is free and publically available on the GitHub deposit: https://github.com/atalha-cmd/Biophysics_ML.git maintained by SMU graduate student Md Abu Talha.
-It contains the Fortran code for generating the electrostatic features using Cartesian treecode, the Python code to generate topological features (a modified version as used in [19]), and the Python code for the neural network, which is originated from Elyssa Sliheet’s thesis work [66].
+Figure 10: Scatter plots for the prediction of $E_{\text{solv}}$ on Dataset 1: (a) topological features only, (b) e-features only ($p=2$, $L=1$ with total 90 features), and (c) both features ($p=3$, $L=0$ with total 20 e-features).
+3.2 Additional Comparison and Justification
+We here provide some comparison studies between the proposed models with baseline models, surrogate models, as well as different grouping and persist homologies strategies.
+In addition, some details about the cross-validation in the training process from a selected case is provided.
+The limitations of the proposed model and the features generation algorithms are also discussed.
+3.2.1 $E_{\text{coul}}$ predicted from persistence image on Dataset 2
+(a)                 (b)                 (c)
+Figure 11: Prediction performance of $E_{\mathrm{coul}}$ on Dataset 2 using combined features: (a) $R^{2}$ Score, (b) scaled MSE, and (c) scaled MAPE as functions of the number of features, comparing Barcode and PImage representations.
+As explained previously, the topological features can be extracted using either barcodes or persistent images.
+When only the persistence image-based topological features are used, the model achieves an MSE of 0.037, MAPE of 0.082 and an $R^{2}$ value of 0.964, which is comparable to the performance obtained using barcode-based topological features (MSE = 0.030, MAPE = 0.071 and $R^{2}$ = 0.970).
+Both approaches demonstrate excellent predictive accuracy even in the absence of electrostatic features.
+When these topological features are combined with electrostatic features, the barcode-based representation yields slightly better performance than the persistent image–based representation, as shown in Fig.
+11, in terms of $R^{2}$, MSE, and MAPE.
+Considering its relatively better performance, lower memory usage, and more efficient computation, we adopt the barcode-based approach throughout the paper.
+It worths noting, however, that the persistent image approach may offer advantages in terms of stability and vectorized structure, which will be investigated further in future work.
+3.2.2 Robustness of $E_{\text{coul}}$ prediction to homology-aware data splitting (Dataset 2)
+Figure 12: Performance comparison for $E_{\mathrm{coul}}$ prediction on Dataset 2 under random splitting, UniProt-based grouping, and MMseqs sequence-identity clustering (30% identity, 80% coverage), using combined features.
+In our machine learning model, we first randomly divide the dataset into training sets and testing sets with proportions of 80% and 20%, respectively.
+The training set is further divided into five equal subsets for five-fold cross-validation, from which the model with the best validation performance is selected and subsequently evaluated on the held-out test set.
+A potential concern is that the PDBbind dataset contains multiple structures corresponding to the same protein or closely related homologs (e.g., alternative conformations, mutants, or orthologous proteins).
+If such similar proteins are distributed in both training and testing sets, data leakage may artificially inflate performance.
+To mitigate this risk, we adopted two complementary strategies.
+First, we grouped PDB entries according to their UniProt (Universal Protein Resource) accession UniProt:2023, ensuring that all structures of the same protein were assigned entirely to either the training or testing set.
+However, since proteins with different UniProt IDs may still share high sequence similarity, this alone does not eliminate homologous overlap.
+We therefore applied MMseqs (Many-against-Many sequence searching) MMseqs:2017 clustering at 30% sequence identity and 80% coverage.
+Chain sequences were clustered, and PDB entries were grouped by connected components such that any entries sharing sequence similarity, including transitive relationships, were placed in the same group.
+Dataset splitting was then performed at the group level to prevent sequence-similarity leakage.
+As shown in Fig.
+12, similar trends in $R^{2}$, scaled MSE, and scaled MAPE are observed across random, UniProt-based, and MMseqs-based splits.
+The slightly better performance under random splitting is expected due to less stringent generalization.
+Comparable results under these methods indicate that the model’s performance is not significantly driven by the memorization of homologous sequences.
+3.2.3 Cross-Validation Performance
+(a)                        (b)
+Figure 13: Cross-validation training histories for $E_{\mathrm{coul}}$ prediction ($p=4$ and $L=1$) on Dataset 2; (a) Electrostatic features only, (b) Combined features.
+The mean training and validation losses are shown as solid lines, with shaded bands representing standard deviation across folds.
+Now we investigate the training process that uses five-fold cross-validation for over-fitting checks.
+In Fig.
+13, we report the cross-validation training histories or learning curves for $E_{\mathrm{coul}}$ prediction ($p=4$ and $L=1$) on Dataset 2.
+From the mean training and validation loss curves and the corresponding one-standard-deviation bands, we can see that the selected model is stable and free of overfitting.
+Meanwhile, the performance of using both types of features is significantly better than using electrostatic features only, as indicated by faster convergence and narrower bands.
+Similar trends are observed for models predicting other physical quantities across different datasets.
+We further show in Fig.
+14 the mean, 95% confidence interval (assuming $t$-distribution), and one standard deviation band of $R^{2}$ in (a) and MSE in (b) from the validation results of the five-fold cross-validation training procedure across different numbers of electrostatic features when combined features are used.
+The narrow band and interval widths demonstrate the stability of the model.
+Note that for $E_{\mathrm{coul}}$, the advantages of using more electrostatic features are not obvious, since the contribution of topological features dominates.
+Figure 14: Validation performance during training across all cross-validation folds using combined features.
+Solid lines show mean performance across folds, while shaded regions and error bars indicate 95% confidence intervals and $\pm$ one standard deviation, respectively.
+3.2.4 Comparison with baseline models
+Figure 15: Prediction performance of the proposed DNN model compared with baseline methods—Ridge Regression, Random Forest, and Gradient Boosting—across varying numbers of input features.
+We next present the comparison results between the deep neural network model and other baseline models, including ridge regression, random forest, and gradient boosting.
+All methods using the same number of electrostatic and topological features, with the number of electrostatic features increasing as shown in Fig.
+15.
+From the comparison, we observe that DNN, Random Forest, and Gradient Boosting, as more flexible methods, perform significantly better than Ridge Regression.
+The DNN shows slight improvements over Random Forest and Gradient Boosting in terms of $R^{2}$ and MSE.
+3.2.5 CPU time comparison and Size-Dependent Performance
+In order to show the efficiency of the learned model for predicting $E_{\text{solv}}$ compared with the MIBPB solver, we report the CPU time and solvation energies $E_{\text{solv}}$ obtained from both approaches in Fig.
+16.
+As shown in (a), the predicted $E_{\text{solv}}$ values are closely aligned; whereas in (b), the time required by the learned model in red is significantly shorter than the runtime of the MIBPB solver in black, which increases rapidly with molecular size.
+Also as shown in Fig.
+16(b), the total machine learning pipeline time in red is further decomposed into time to compute electrostatic features in blue, time to compute topological features in orange, the time to predict solvation energies using the learned model in green.
+(a)                        (b)
+Figure 16: Performance comparison between the proposed DNN model and the MIBPB solver.
+(a) $E_{\text{solv}}$ predicted by the learned model versus that is computed by MIBPB solver selected protein structures; (b) Computational runtime with respect to protein size (number of atoms), demonstrating the significant efficiency of the DNN model.
+The MIBPB solver is evaluated with grid size $h=0.5$, dielectric constants $\epsilon_{1}=1$, $\epsilon_{2}=80$, and inverse Debye length $\kappa=0$.
+To further evaluate the robustness of the model with respect to protein size, Fig.
+17 presents a size-dependent performance analysis.
+The scatter plot (predicted value vs reference value) in (a) and Box-and-whisker plot (absolute percentage error) in (b) both indicate the model’s uniform performance across protein sizes.
+(a)                        (b)
+Figure 17: Protein size–dependent performance of the DNN model.
+(a) Predicted versus reference $E_{\mathrm{solv}}$ colored by protein size.
+(b) Distribution of absolute percentage error across protein size ranges.
+3.2.6 Model comparison with testing hypothesis
+Theoretically, the three research purposes as mentioned earlier can be treated as statistical hypotheses and evaluated through testing hypothesis for model comparison.
+However, this would require pairwise comparisons between models with numerous simulations to generate statistically significant metrics, which is a very costly procedure.
+Here, we present one such case to illustrate the approach.
+We select a particular set of electrostatic features $(p=2,L=1)$ to compare three models $M_{e}$, $M_{t}$, and $M_{e,c}$, which use electrostatic features only, topological features only, and combined features, respectively.
+These three models are used to predict $E_{\text{solv}}$ on Dataset 2 using 10 independent random 80–20 train–test splits, and the corresponding $R^{2}$ and MSE values are recorded.
+The resulting $R^{2}$ and MSE metrics, summarized in Table 4, are then used to test the following hypotheses:
+$H_{0}$: $M_{e,t}$ performs as well as $M_{e}$ or $M_{t}$
+$H_{a}$: $M_{e,t}$ performs better than $M_{e}$ or $M_{t}$
+Table 4: Metrics obtained by running models $M_{e}$, $M_{t}$, and $M_{e,t}$ for the prediction of $E_{\text{solv}}$ using 10 independent random 80–20 train–test splits on Dataset 2.
+Metrics $S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$ $S_{5}$ $S_{6}$ $S_{7}$ $S_{8}$ $S_{9}$ $S_{10}$ $\bar{S_{i}}$
+$R^{2}(M_{e})$ 0.876 0.885 0.883 0.853 0.876 0.859 0.879 0.862 0.873 0.885 0.873
+$R^{2}(M_{t})$ 0.838 0.802 0.816 0.847 0.857 0.803 0.844 0.812 0.852 0.810 0.828
+$R^{2}(M_{e,t})$ 0.935 0.936 0.907 0.937 0.926 0.887 0.931 0.914 0.921 0.899 0.919
+[…]
+[…]
+[…]
+The four paired one-sided Wilcoxon signed-rank tests on $R^{2}$ and MSE between $M_{e,t}$ and $M_{e}$ or $M_{t}$ all return $p$-value < 0.001 thus the null hypotheses are rejected, indicating model $M_{e,t}$ is significantly better than $M_{e}$ or $M_{t}$.
+3.2.7 Limitation
+The manuscript focuses on providing algorithms for feature generation for proteins in general and is not restricted by protein size or type.
+However, we do notice some limitations for the proposed methods, as listed below.
+(1) The electrostatic features primarily focus on the protein electrostatic profile therefore they have limited impact on capturing the interplay between the protein and its surrounding environment.
+This limitation is reflected in the simulation results, where Coulombic energy prediction performs better than solvation energy prediction.
+In future work, we plan to enhance the electrostatic features by incorporating reaction potentials generated from the GB model at atomic centers to better characterize protein-solvent interactions.
+(2) When the dataset size is small, for the same types of features, the advantages of DNN over Gradient Boosting or Random Forrest are difficult to justify.
+(3) Identifying the optimal values of $p$ and $L$ for generating electrostatic features is challenging, as they depend on factors such as dataset size, response types, and the performance of topological features.
+This limitation requires multiple tests to determine the optimal $p$ and $L$ during model training.
+We are planning to implement and test the Barycentric treecode approach Wang:2020 as explained in the supplementary material for potential solutions.
 4 Conclusion
-The project provides a novel approach to generate topological and electrostatic features combined to represent protein structures and charge distributions in a physical-informed multiscale and uniform way.
-Using these features, we can predict important protein properties such as the solvation energy and Coulombic energies.
-The numerical simulation validates our three purposes such that larger dataset, more electrostatic features, and combined features can improve the performance of the machine learning model.
-Other than predicting solvation energies and Coulomb energy, the resulting electrostatic and topological features potentially have broad usage as far as the proteins are needed to be represented in the forms of uniform and multiscale features for machine learning models.
-In the future, we will further investigate the role of reaction potential, which will be calculated using the much faster GB model as explained in the manuscript.
-We will also investigate Barycentric Lagrange treecode to generate electrostatic features as compared with the Cartesian treecode.
+The project provides a novel approach for generating combined topological and electrostatic features to represent protein structures and charge distributions in a physical-informed multiscale and uniform manner.
+Using these features, we can predict important protein properties such as solvation energy and Coulomb energy.
+The numerical simulations validate our three main findings: larger dataset, higher-resolution electrostatic features, and the combination of features all improve the performance of the ML model.
+Beyond predicting solvation and Coulomb energies, the resulting electrostatic and topological features have the potential for broad applicability, particularly in representing proteins as uniform and multiscale features for ML models.
+In the future work, we will further investigate the role of the reaction potential, which will be computed using the much faster GB model, as discussed in the manuscript.
+We will also explore the use of the barycentric Lagrange treecode for generating electrostatic features, in comparison with the Cartesian treecode.
+5 Data and Software Availability
+The code and data associated with this project are freely and publicly available in the GitHub repository: https://github.com/atalha-cmd/Biophysics_DNN, maintained by SMU graduate student Md Abu Talha.
+It contains the Fortran code for generating electrostatic features using Cartesian treecode, the Python code to generate topological features (a modified version as used in Cang:2017), the Python code for the neural network originating from Elyssa Sliheet’s thesis work Sliheet:2024, and the data required to reproduce the reported results.
+{acknowledgement}
+This work of ES, MAT, and WG was supported in part by National Science Foundation (NSF) grants DMS-2110922 and DMS-2110869.
+ES was also partially supported by the NSF RTG-1840260 grant.
+We thank the SMU O’Donnell Data Science and Research Computing Institute (ODSRCI) for providing computing hardware and software.
+These resources together made this project possible.
+{suppinfo}
+The following files are available free of charge.
+  • supple-achemso.pdf: this file contains additional information of models, methods, and results which are less essential but could assist the readers in better understanding or reproducing the proposed work.